### v1

# Injectives obstruct Fourier-Mukai functors

Abstract.
We use the inclusion of injectives into the canonical heart as a replacement for tilting objects in computations of the characteristic morphism. Then we apply this construction to proofs of non-liftability of candidate non-Fourier-Mukai functors, i.e. functors that do not admit an $\mathcal{A}_{\infty}$/$\mathrm{dg}$-lift. This approach allows explicit computation of the obstruction against an $\mathcal{A}_{\infty}$-lift. We in particular observe that this computation gives for smooth degree $d>2$ hypersurfaces an abundance of non-Fourier-Mukai functors.
footnotetext: The author is a Postdoctoral Research Fellow at the Université Libre de Bruxelles. The research in this work has been supported by the MIS (MIS/BEJ - F.4545.21) Grant from the FNRS, a mobility grant by the FNRS and an ACR Grant from the Université Libre de Bruxelles.
1. Introduction
The idea behind Fourier-Mukai functors as categorified correspondences was introduced by S. Mukai [Muk81] to understand the relation between the derived category of an abelian variety $\mathcal{A}$ and its dual $\mathcal{A}^{\vee}$. Its classical definition being:
Definition 1.1.
[Huy06] A functor
$f:\mathcal{D}^{b}\left(\mathrm{coh}X\right)\to\mathcal{D}^{b}\left(\mathrm{coh}Y\right),$
respectively
$f\colon\mathcal{D}^{b}\left(\mathrm{Qcoh}X\right)\to\mathcal{D}^{b}\left(\mathrm{Qcoh}Y\right),$
for two schemes $X$ and $Y$ is Fourier-Mukai if there exists an $M$ in $\mathcal{D}^{b}\left(\mathrm{coh}\left(X\times Y\right)\right)$, respectively in $\mathcal{D}^{b}\left(\mathrm{Qcoh}\left(X\times Y\right)\right)$, such that
$f\cong\pi_{1,*}\left(M\otimes\pi_{2}^{*}\_\right).$
This concept allows the study of functors from a geometric point of view and in particular the application of geometric arguments and constructions to them. With time it was realized that in fact every known functor between derived categories of smooth projective schemes is Fourier-Mukai. For example push-forward and pull-back along a morphism of schemes $f$ is induced by taking $M$ to be the structure sheaf of its graph $\mathcal{O}_{\Gamma_{f}}$.
A celebrated result by D. Orlov gave a criterion to check whether a functor is Fourier-Mukai,
Theorem 1.2.
[Orl97] Let $f:\mathcal{D}^{b}\left(\mathrm{coh}X\right)\to\mathcal{D}^{b}\left(\mathrm{coh}Y\right)$ be a fully-faithful functor between derived categories of smooth projective schemes. Then $f$ is Fourier-Mukai.
Whereas B. Toën realized a beautiful conceptual categorization of the property,
Theorem 1.3.
[Toe07] Let $f:\mathcal{D}^{b}\left(\mathrm{Qcoh}X\right)\to\mathcal{D}^{b}\left(\mathrm{Qcoh}Y\right)$ be an exact functor between derived categories of smooth projective schemes. Then $f$ is Fourier-Mukai if and only if there exists a $\mathrm{dg}$-lift of $f$ i.e.
$\widehat{f}:\mathcal{D}_{\mathrm{dg}}^{b}\left(\mathrm{Qcoh}X\right)\to\mathcal{D}_{\mathrm{dg}}^{b}\left(\mathrm{Qcoh}Y\right),$
where $\mathcal{D}_{\mathrm{dg}}^{b}\left(\_\right)$ is a $\mathrm{dg}$-enhancement of $\mathcal{D}^{b}\left(\_\right)$.
This Theorem in particular allows the interpretation of Fourier-Mukai functors as the geometric analog to bimodules and their represented functors. These observations led to a conjecture that actually all exact functors between derived categories of smooth projective schemes are Fourier-Mukai.
Conjecture 1.4.
[Lun07] Let $f:\mathcal{D}^{b}\left(\mathrm{coh}X\right)\to\mathcal{D}^{b}\left(\mathrm{coh}Y\right)$ be a functor between derived categories of smooth projective schemes. Then $f$ is Fourier-Mukai.
More than ten years later A. Rizzardo, M. Van den Bergh and A. Neeman constructed the first non-Fourier-Mukai functor.
Theorem 1.5.
[RVdBN19] There exists an exact functor
$\psi_{\eta}:\mathcal{D}^{b}\left(\mathrm{coh}Q_{3}\right)\to\mathcal{D}^{b}\left(\mathrm{coh}\mathbb{P}^{4}\right)$
that is not Fourier-Mukai, where $Q_{3}\hookrightarrow\mathbb{P}^{4}$ is a smooth three dimensional quadric.
We were able to generalize their argument to higher dimensions in [Kü22] using twisted Hodge diamonds, while T. Raedschaelders, A. Rizzardo and M. Van den Bergh [RRVdB19] were able to construct similar non-Fourier-Mukai functors by losing control over the target space.
A big drawback of all these approaches is that they heavily rely on the existence of tilting objects in the source category in order to control liftability. We avoid this issue here by replacing the tilting bundle with the subcategory of injectives in the canonical heart of the derived category of quasi-coherent sheaves. This observation allows the computation of the equivariant characteristic morphism with respect to the category of injectives and explicit computations of the induced $\mathcal{A}_{\infty}$-obstructions. In particular we are able to prove the following very general result:
Theorem 8.11.
Let $f:X\hookrightarrow Y$ be a divisor and let $0\neq\eta\in\mathrm{HH}^{*}\left(X,M\right)$ be a Hochschild cohomology class such that $f_{*}\eta=0$ for $M$ a linebundle on $X$. Then there exists a non-Fourier-Mukai functor
$\psi_{\eta}:\mathcal{D}^{b}\left(\mathrm{Qcoh}X\right)\to\mathcal{D}^{b}\left(\mathrm{Qcoh}Y\right).$
Explicit computations for smooth hypersurfaces in projective space then yield:
Corollary 8.12.
Let $f:X\hookrightarrow\mathbb{P}^{n+1}$ be a smooth degree $d$-hypersurface. Then the space of non-Fourier-Mukai functors arising as deformation of the shape $\psi_{\eta}$ can be computed explicitly.
1.1. Proof strategy
By classical methods $\mathcal{A}_{\infty}$-categories are equivalent to $\mathrm{dg}$-categories and better behaved with respect to liftability. So we will work with $\mathcal{A}_{\infty}$-structures in order to obstruct $\mathrm{dg}$/$\mathcal{A}_{\infty}$-lifts.
Let us briefly sketch the above mentioned previous approaches to non-liftability. These assumed the existence of a tilting bundle $T$ with endomorphism algebra $\Gamma$. After applying the candidate non-Fourier-Mukai functor $\psi_{\eta}$ this turns $\psi_{\eta}\left(T\right)$ into a $\Gamma$-equivariant object in $\mathcal{D}^{b}\left(Y\right)$, whose $\mathcal{A}_{\infty}$-liftability then can be obstructed.
The key insight in our approach is that the non-liftability of $\psi_{\eta}\left(T\right)$ boils down to the the functor $\psi_{\eta}\circ\iota_{T}:*_{\Gamma}\to\mathcal{D}^{b}\left(Y\right)$ not admitting an $\mathcal{A}_{\infty}$-lift, here we denoted by $*_{\Gamma}$ the $𝕜\Bbbk$-linear category with one object * and endomorphisms $\Gamma$. In particular, since $T$ was assumed to be tilting the functor
$\iota_{T}:*_{\Gamma}\to\mathcal{D}^{b}\left(X\right)$
determines functors on $\mathcal{D}^{b}\left(X\right)$ uniquely.
Our approach consists of replacing the functor $\iota_{T}:*_{\Gamma}\hookrightarrow\mathcal{D}^{b}\left(X\right)$ with the functor given by the inclusion of the full subcategory of injective objects into the heart
$\iota:\mathrm{Inj}\hookrightarrow\mathcal{D}^{b}\left(\mathrm{Qcoh}X\right).$
By classical homological algebra this functor completely controls $\mathcal{D}^{b}\left(\mathrm{Qcoh}X\right)$ [Wei94].
In order to explain our proof strategry, consider the follwoing diagram
$\mathrm{Inj}$$\mathcal{D}^{b}\left(\mathrm{Qcoh}X\right)$$\mathcal{D}^{b}\left(\mathcal{X}_{\eta}\right)$$\mathcal{D}^{b}\left(\mathrm{Qcoh}Y\right),$$\mathcal{D}_{\mathrm{dg}}^{b}\left(\mathrm{Qcoh}X\right)$$\mathcal{D}_{\mathrm{dg}}^{b}\left(\mathcal{X}_{\eta}\right)$$\mathcal{D}_{\mathrm{dg}}^{b}\left(\mathrm{Qcoh}Y\right)$.$\iota$$\iota$$L$$\widetilde{f}_{*}$$\widetilde{f}_{*}$
where to top row indicates the $\mathrm{dg}$-enhancements of the bottom row. The functor $\psi_{\eta}$ will be constructed as the composition of $L$ and $\widetilde{f}_{*}$.
As $\iota$ uniquely determines $\mathrm{dg}$-functors from the category $\mathcal{D}^{b}\left(\mathrm{Qcoh}\left(X\right)\right)$ we can compute its equivariant characteristic morphism in order to obstruct an $\mathcal{A}_{\infty}$-lift of $L$. This yields a non-vanishing $\mathcal{A}_{\infty}$-obstruction $o_{m}\left(L\circ\iota\right)$. The plan is to push this obstruction forward. In contrast to [RVdBN19] and [Kü22] this needs to be done with a lot of care as we do not have access to Serre duality in our setting. In particular we study the concrete construction of $\widetilde{f}_{*}$ to compute that the obstruction $o_{m}\left(\psi_{\eta}\circ\iota\right)$ indeed is independent of choices and coincides with the pushed forward obstruction $f_{*}o_{m}\left(L\circ\iota\right)$. This means that there cannot be an $\mathcal{A}_{\infty}$-lift of $\psi_{\eta}\circ\iota$. In particular, as $\iota$ is the inclusion of injectives, it is itself a $\mathrm{dg}$-functor and hence we conclude that $\psi_{\eta}$ cannot have an $\mathcal{A}_{\infty}$-lift as desired.
1.2. Structure of the paper
In Section 2.1 we recall preliminaries on Hochschild coholomology and the characteristic morphism in the classical setting. In particular we recall in Definition 2.5 the construction of an infinitesimal deformation along a Hochschild cohomology class, which is a key player in the construction of $\psi_{\eta}.$
Then we recall in Section 2.2 how one can interpret equivariant objects as functors out of one object categories. This will be used later to make our arguments on the characteristic morphism of subcategories clearer.
Using these observations, together with insights from [Toe07], in Section 3 we compute that the $\mathrm{Inj}$-equivariant characteristic morphism of the subcategory $\mathrm{Inj}$ does not vanish. This construction will be our replacement for the assumption on the existence of a tilting object.
We recall in Section 4 the construction from [RVdBN19] of candidate non-Fourier-Mukai functors. In Section 5 we then collect the properties, definitions and computations of $\mathcal{A}_{\infty}$-obstructions that we will use later.
Since we are not using tilting objects in the category of coherent sheaves we are not able to use Serre duality as in [RVdBN19, Kü22] in order to show that the relevant $\mathrm{Ext}$-groups are concentrated in degree $-1$. So we will have to control the $\mathcal{A}_{\infty}$-structure on our target category more explicitly. For this we study in Section 6 the explicit structure of the functor $\widetilde{f}$. In particular we completely understand the essential image and its $\mathcal{A}_{\infty}$-bimodule structure.
Then we finish in Section 7 by computing explicitly that the obstructions of $\psi_{\eta}$ cannot vanish. We then combine this result with computations from [Kü22] in order to conclude the existence of a vast amount of non-Fourier-Mukai functors with well-behaved target categories.
Acknowledgements
I would like to thank the SL-Math institute for hosting the semester on “non-commutative algebraic geometry” where a lot of this work has been completed. Furthermore I want to show my gratitude to M. Van den Bergh and A. Rizzardo for helpful discussions and deep insights and Š. Špenko for her awesome support and helpful discussions. Finally I would like to thank B. Briggs and A. Cǎldǎraru for their insights on Hochschild cohomology, $\mathcal{A}_{\infty}$-structures and formality of sheaves over source spaces.
Notation
Let $𝕜\Bbbk$ denote a field of characteristic zero. We consider all schemes and categories to be over $𝕜\Bbbk$. For a $𝕜\Bbbk$-linear category $\mathcal{X}$, denote by $\mathcal{D}^{b}\left(\mathcal{X}\right)$ the bounded derived category of modules over $\mathcal{X}$. We will consider left modules. Denote by $\mathcal{D}_{\mathrm{dg}}^{b}\left(\_\right)$ the (canonical) dg-enhancement of $\mathcal{D}^{b}\left(\_\right)$. We denote by $\mathcal{D}^{b}_{\mathcal{C}}\left(\_\right)$ the full subcategory of $\mathcal{D}^{b}\left(\_\right)$ with cohomology objects in $\mathcal{C}$. We will also refer to $\mathrm{dg}$ respectively $\mathcal{A}_{\infty}$-functors between smooth geometric categories as Fourier Mukai functors following the insight from B. Toën [Toe07]. We also use the shorthand
$\left[f_{i},m_{k}\right]:=$ $\sum_{\mathclap{r+t=i-1}}\left(-1\right)^{r+kt}f_{i}\left(\mathrm{Id}^{\otimes r}\otimes\mathrm{m}_{k}\otimes\mathrm{Id}^{\otimes t}\right)$
$-\sum_{\mathclap{1\leq j\leq k}}\left(-1\right)^{ij}\mathrm{m}_{k}\left(\alpha_{1}\otimes...\otimes f_{i}\left(\alpha_{j+1},...,\alpha_{j+i}\right)\otimes...\otimes\alpha_{k+i}\right)$
for a morphism $\mathcal{A}^{\otimes i}\to\mathcal{M}$ and $\mathcal{M}$ an $\mathcal{A}_{\infty}$-$\mathcal{A}$-module.
2. Preliminaries
We start off by recalling some basic notions needed for our constructions and arguments. For the basic notions on $\mathcal{A}_{\infty}$-structures we refer to the notes by B. Keller [Kel01], the work of H. Lefevre [LH03] and P. Seidel [Sei08]. For the construction of $\mathrm{dg}$-hulls we refer to [KKS09].
2.1. Preliminaries Hochschild cohomology and characteristic morphism
As we will work a lot with obstructions and Hochschild cohomology we recall the basic constructions and notions here. For a more in depth discussion on the incarnations of Hochschild cohomology we use here, we refer the reader to the original work of G. Hochschild [Hoc45], R. Buchweitz and H. Flenner [BF08] respectively R. Swan [Swa96] for its geometric incarnation, and for its relation to deformation theory to the work of M. Van den Bergh and W. Lowen [LVdB05]. Finally we want to refer for a treatment on the characteristic morphism to the work of B. Briggs and V. Gelinas [BG17].
Definition 2.1.
Let $\mathcal{X}$ be a small $𝕜\Bbbk$-linear category and $\mathcal{M}$ a $𝕜\Bbbk$-central $\mathcal{X}$-bimodule. Then the Hochschild complex of $\mathcal{X}$ with values in $\mathcal{M}$ is given by the complex
[…]
with differential given by
$df\left(x_{1}\otimes\cdot\otimes x_{n+1}\right):=$ $x_{1}f\left(x_{2}\otimes\cdots\otimes x_{n+1}\right)$
$+\sum_{i=1}^{n}\left(-1\right)^{i}f\left(x_{1}\otimes\cdots\otimes x_{1}x_{i+1}\otimes\cdots\otimes x_{n+1}\right)$
$+\left(-1\right)^{n+1}f\left(x_{1}\otimes\cdots\otimes x_{n}\right)x_{n+1}.$
The Hochschild cohomology $\mathrm{HH}^{*}\left(\mathcal{X},\mathcal{M}\right)$ is the cohomology of $\mathrm{CC}^{*}\left(\mathcal{X},\mathcal{M}\right)$.
The following classical Lemma 2.2 allows for a more streamlined construction and will help us interpret the Hochschild cohomology in more generality later. In particular it highlights the importance of $\mathcal{M}$ being a bimodule.
Lemma 2.2.
Let $\mathcal{X}$ be a $𝕜\Bbbk$-linear category and let $\mathcal{M}$ be a $𝕜\Bbbk$-central bimodule over $\mathcal{X}$. Then the Hochschild cohomology of $\mathcal{X}$ with values in $\mathcal{M}$ is
$\mathrm{HH}^{*}\left(\mathcal{X},\mathcal{M}\right):=\mathrm{Ext}^{*}_{\mathcal{X}\otimes\mathcal{X}^{\mathrm{op}}}\left(\mathcal{X},\mathcal{M}\right).$
Definition 2.3.
[…]
$\left(x,i\right)\in\mathrm{obj}\left(\mathcal{X}\right)\times\mathrm{obj}\left(\mathcal{I}\right),$
morphism spaces given by
[…]
and higher compositions
[…]
Remark 2.4.
In bigger generality constructing tensor products of $\mathcal{A}_{\infty}$-categories needs to be done with a lot of care [LRG20]. However, we will only consider tensor products with $𝕜\Bbbk$-linear categories, which are a lot better behaved and allow our naive approach to work out.
The restriction in the following definition to $m\geq 3$ is not necessary in general. However, as we will have naturally this assumption and it makes the formulas a lot more concise we present it in this special case.
Definition 2.5.
Let $\mathcal{X}$ be a $𝕜\Bbbk$-linear category and let $\mathcal{M}$ be a $𝕜\Bbbk$-central $\mathcal{X}$-bimodule. Let $\eta\in\mathrm{CC}^{m}\left(\mathcal{X},\mathcal{M}\right)$ be a cocycle with $m\geq 3$. Then the infinitesimal deformation of $\mathcal{X}$ along $\eta$ is the $\mathcal{A}_{\infty}$-category $\mathcal{X}_{\eta}$ given by objects $\mathrm{obj}\mathcal{X}_{\eta}=\mathrm{obj}\mathcal{X}$, morphisms
$\mathcal{X}_{\eta}\left(x,y\right):=\mathcal{X}\left(x,y\right)\oplus\mathcal{M}\left(x,y\right)\left[2-m\right]$
and actions given by
$\mathrm{m}_{2}:\left(\mathcal{X}\left(x,y\right)\oplus\mathcal{M}\left(x,y\right)\left[2-m\right]\right)^{\otimes 2}$ $\to\mathcal{X}\left(x,y\right)\oplus\mathcal{M}\left(x,y\right)\left[2-m\right]$
$\left(x,m\right)\otimes\left(x^{\prime},m^{\prime}\right)$ $\mapsto\left(xx^{\prime},xm+m^{\prime}x\right)$
$\mathrm{m}_{m}:\left(\mathcal{X}\left(x,y\right)\oplus\mathcal{M}\left(x,y\right)\left[2-m\right]\right)^{\otimes m}$ $\to\mathcal{X}\left(x,y\right)\oplus\mathcal{M}\left(x,y\right)\left[2-m\right]$
$x_{1}\otimes...\otimes x_{m}$ $\mapsto\eta\left(x_{1},...,x_{m}\right)$
This construction comes with a canonical morphism that is the identity on $\mathcal{X}$ and zero on $\mathcal{M}$,
(1) $\mathcal{X}_{\eta}\to\mathcal{X}.$
Proposition 2.6.
[RVdBN19, Lemma 6.1.1.] Let $\mathcal{X}$ be a $𝕜\Bbbk$-linear category, $\mathcal{M}$ a $𝕜\Bbbk$-central $\mathcal{X}$-bimodule and $\eta,\mu\in\mathrm{CC}^{m}\left(\mathcal{X},\mathcal{M}\right)$ such that $\eta=\mu$ in $\mathrm{HH}^{*}\left(\mathcal{X},\mathcal{M}\right)$. Then
$\mathcal{X}_{\eta}\cong\mathcal{X}_{\mu}.$
Remark 2.7.
Let $\mathcal{X}$ and $\mathcal{I}$ be $𝕜\Bbbk$-linear categories and $\mathcal{M}$ a $𝕜\Bbbk$-central bimodule. Then there is a morphism of Hochschild complexes
$\mathrm{CC}^{*}\left(\mathcal{X},\mathcal{M}\right)$ $\to\mathrm{CC}^{*}\left(\mathcal{X}\otimes\mathcal{I},\mathcal{M}\otimes\mathcal{I}\right)$
$\eta$ $\mapsto\eta\cup\mathrm{Id}$
One can compute that this morphism is compatible with deformations, i.e.
$\mathcal{X}_{\eta}\otimes\mathcal{I}\cong\left(\mathcal{X}\otimes\mathcal{I}\right)_{\eta\cup\mathrm{Id}}.$
Definition 2.8.
Let $\mathcal{X}$ be a $𝕜\Bbbk$-linear category, $\mathcal{M}$ a $𝕜\Bbbk$-central bimodule and let $\mathcal{G}$ be a $\mathcal{X}$-module. Then _$\_\otimes\mathcal{G}$ induces a map called the characteristic morphism
$c_{\mathcal{G}}:\mathrm{HH}^{*}\left(\mathcal{X},\mathcal{M}\right)\cong\mathrm{Ext}^{*}_{\mathcal{X}\otimes\mathcal{X}^{\mathrm{op}}}\left(\mathcal{X},\mathcal{M}\right)\xrightarrow{\otimes_{\mathcal{X}}\mathcal{G}}\mathrm{Ext}^{*}_{\mathcal{X}}\left(\mathcal{G},\mathcal{M}\otimes\mathcal{G}\right).$
Analogously we have the dual characteristic morphism
$c^{*}_{\mathcal{G}}:\mathrm{HH}^{*}\left(\mathcal{X},\mathcal{M}\right)\cong\mathrm{Ext}^{*}_{\mathcal{X}\otimes\mathcal{X}^{\mathrm{op}}}\left(\mathcal{X},\mathcal{M}\right)\xrightarrow{\mathrm{RHom}_{\mathcal{X}}\left(\_,\mathcal{G}\right)}\mathrm{Ext}^{*}_{\mathcal{X}}\left(\mathrm{RHom}\left(\mathcal{M},\mathcal{G}\right),\mathcal{G}\right),$
induced by $\mathrm{RHom}_{X}\left(\_,\mathcal{G}\right)$.
2.2. Equivariant objects over categories
In this section we recall the basic notions of equivariant objects and how to interpret them as functors from a small category, which we will apply in the next section to compute the equivariant characteristic morphisms of these objects.
We start by recalling the basic definition.
Definition 2.9.
Let $\mathcal{C}$ be a $𝕜\Bbbk$-linear category and let $\Gamma$ be a $𝕜\Bbbk$-algebra. A $\Gamma$-equivariant object $X\in\mathcal{C}$ is an object equipped with an action morphism of $𝕜\Bbbk$-algebras
$\varphi:\Gamma\to\mathcal{C}\left(X,X\right).$
A morphism between $\Gamma$-equivariant objects $X$ and $Y$ is a morphism $f:X\to Y$ in $\mathcal{C}$ such that
$\varphi_{Y}\left(\gamma\right)\circ f=f\circ\varphi_{X}\left(\gamma\right).$
There also exists a $\Gamma$-equivariant incarnation of the characteristic morphism, given by:
Definition 2.10.
Let $\mathcal{X}$ be a $𝕜\Bbbk$-linear category, $\mathcal{M}$ a $𝕜\Bbbk$-central bimodule and let $\mathcal{G}$ be a $\Gamma$-equivariant $\mathcal{X}$-module. Then _$\_\otimes\mathcal{G}$ induces the the $\Gamma$-equivariant characteristic morphism
$c_{\Gamma,\mathcal{G}}:\mathrm{HH}^{*}\left(\mathcal{X},\mathcal{M}\right)\to\mathrm{Ext}^{*}_{\mathcal{X}\otimes\Gamma^{\mathrm{op}}}\left(\mathcal{G},\mathcal{M}\otimes\mathcal{G}\right).$
While $\mathrm{RHom}\left(\_,\mathcal{G}\right)$ induces the the $\Gamma$-equivariant dual characteristic morphism
$c^{*}_{\Gamma,\mathcal{G}}:\mathrm{HH}^{*}\left(\mathcal{X},\mathcal{M}\right)\to\mathrm{Ext}^{*}_{\mathcal{X}\otimes\Gamma^{\mathrm{op}}}\left(\mathrm{RHom}\left(\mathcal{M},\mathcal{G}\right),\mathcal{G}\right).$
Lemma 2.11.
Let $\mathcal{C}$ be a $𝕜\Bbbk$-linear category and let $\Gamma$ be a $𝕜\Bbbk$-algebra. Then a $\Gamma$-equivariant object $X$ is equivalently a functor from the one object category $*_{\Gamma}$ with endomorphisms $\Gamma$,
$\iota_{X,\Gamma}:*_{\Gamma}$ $\to\mathcal{C}$
* $\mapsto X.$
Furthermore this naturally extends to an equivalence of categories between $\Gamma$-equivariant objects in $\mathcal{C}$ and the functor category $\mathrm{Fun}\left(*_{\Gamma},\mathcal{C}\right)$.
Proof.
Let $\iota_{X,\Gamma}$ be a functor of the given shape, then it defines a $\Gamma$-equivariant object by taking $X=\iota_{X,\Gamma}\left(*\right)$ with action given by
$\iota_{X,\Gamma}:\mathrm{End}\left(*\right)=\Gamma\longrightarrow X=\iota_{X,\Gamma}.$
We can do the same construction the other way around, i.e. make a $\Gamma$-equivariant object $X$ into a functor via fixing the image of * to be $X$ and the structure morphism the map on morphism spaces.
Finally, the naturality requirement for a morphism $\alpha:\iota_{X,\Gamma}\to\iota_{Y,\Gamma}$ in the functor category boils down to
$\alpha\circ\iota_{X,\Gamma}\left(\gamma\right)=\iota_{Y,\Gamma}\left(\gamma\right)\circ\alpha,$
which is precisely the identity from Definition 2.9. So we may identify the two categories. ∎
In particular we may consider equivariant object over small $𝕜\Bbbk$-linear categories by considering functors.
Definition 2.12.
Let $\mathcal{C}$ be a $𝕜\Bbbk$-linear category and $\mathcal{I}$ a small $𝕜\Bbbk$-linear category. An $\mathcal{I}$-equivariant object in $\mathcal{C}$ is a $𝕜\Bbbk$-linear functor
$\mathcal{I}\to\mathcal{C}.$
We use the following notation for the category of equivariant objects over $I$ in $\mathcal{C}$ by
[…]
3. Equivariant characteristic morphism of Injectives
We study the characteristic morphism introduced in the previous chapter for the subcategory of injectives in a category with enough injectives. We start by discussing how we can reinterpret the characteristic morphism by the work of [Toe07]. We then use this insight to prove that we can treat the full subcategory of injectives similarly to a tilting object.
3.1. Interpreting characteristic morphisms of a subcategory as pullback
We use observations from [Toe07] to study in more depth the characteristic morphism
$c_{G}:\mathrm{HH}^{*}\left(\mathcal{C},\mathcal{M}\right)\to\mathrm{Ext}_{\mathcal{C}}^{*}\left(G,\mathcal{M}\otimes_{\mathcal{C}}G\right)$
and the equivariant characteristic morphism
$c_{\Gamma,G}:\mathrm{HH}^{*}\left(\mathcal{C},\mathcal{M}\right)\to\mathrm{Ext}_{\mathcal{C}_{\Gamma}}^{*}\left(G,\mathcal{M}\otimes_{\mathcal{C}}G\right).$
We start by recalling from the previous Section 2.2 that an object $G\in\mathcal{C}$ is equivalently a functor from the one object category $𝕜*_{\Bbbk}$ with endomorphisms $𝕜\Bbbk$ into $\mathcal{C}$. In particular we may consider for an object $G\in\mathcal{C}$ the functor
[…]
respectively for a $\Gamma$-equivariant $T\in\mathcal{C}$ the functor from the one object category $*_{\Gamma}$ with endomorphism $\Gamma$ into $\mathcal{C}$. So we may consider in this case
$\iota_{T}:*_{\Gamma}\to\mathcal{C}.$
Definition 3.1.
Let $\mathcal{C}$ be an abelian category and $I\subset\mathcal{C}$ a subcategory. Then denote by $\iota_{I}:I\hookrightarrow\mathcal{C}\subset\mathcal{D}_{\mathrm{dg}}^{b}\left(\mathcal{C}\right)$ the inclusion of objects of $I$.
Remark 3.2.
By [Toe07] we can interpret the Hochschlid cohomology of a $𝕜\Bbbk$-linear category $\mathcal{C}$ with values in a $𝕜\Bbbk$-central bimodule $\mathcal{M}$ as the following $2$-morphism space in the homotopy category of $\mathrm{dg}$-categories.
$\mathrm{HH}^{*}\left(\mathcal{C},\mathcal{M}\right)\cong\mathrm{Ho}\left(\mathrm{dg}-\mathrm{Cat}\right)\left(\mathcal{D}^{b}_{\mathrm{dg}}\left(\mathcal{C}\right),\mathcal{D}^{b}_{\mathrm{dg}}\left(\mathcal{C}\right)\right)\left(\mathrm{Id},\psi_{\mathcal{M}}\right)$
where $\psi_{\mathcal{M}}$ denotes the $\mathrm{dg}$-functor reperesented by $\mathcal{M}$ i.e.
$\psi_{\mathcal{M}}:\mathcal{D}^{b}\left(\mathcal{C}\right)$ $\to\mathcal{D}^{b}\left(\mathcal{C}\right)$
$X$ $\mapsto\mathcal{M}\otimes X.$
In this interpretation, for a $\Gamma$-equivariant object $T$ the equivariant characteristic morphism turns into application of the pullback functor along the inclusion of the one object category $\iota_{T}:\Gamma\to\mathcal{D}_{\mathrm{dg}}^{b}\left(\mathrm{Qcoh}X\right)$:
$\mathrm{HH}^{*}\left(\mathcal{C},\mathcal{M}\right)\cong\mathrm{Ho}\left(\mathrm{dg}-\mathrm{Cat}\right)\left(\mathcal{D}^{b}_{\mathrm{dg}}\left(\mathcal{C}\right),\mathcal{D}^{b}_{\mathrm{dg}}\left(\mathcal{C}\right)\right)\left(\mathrm{Id},\psi_{\mathcal{M}}\right)$$\mathrm{Ho}\left(\mathrm{dg}-\mathrm{Cat}\right)\left(\Gamma,\mathcal{D}^{b}_{\mathrm{dg}}\left(\mathcal{C}\right)\right)\left(\iota_{\mathcal{G}},\psi_{\mathcal{M}}\circ\iota_{\mathcal{G}}\right)\cong\mathrm{Ext}_{\mathcal{C}_{\Gamma}}\left(\mathcal{G},\mathcal{G}\otimes\mathcal{M}\right).$$c_{\Gamma,\mathcal{G}}\cong\iota_{\mathcal{G},*}$
In particular we can use this interpretation to compute the equivariant characteristic morphism of the subcategory $I\in\mathcal{C}$:
$\mathrm{HH}^{*}\left(\mathcal{C},\mathcal{M}\right)\cong\mathrm{Ho}\left(\mathrm{dg}-\mathrm{Cat}\right)\left(\mathcal{D}^{b}_{\mathrm{dg}}\left(\mathcal{C}\right),\mathcal{D}^{b}_{\mathrm{dg}}\left(\mathcal{C}\right)\right)\left(\mathrm{Id},\psi_{\mathcal{M}}\right)$$\mathrm{Ho}\left(\mathrm{dg}-\mathrm{Cat}\right)\left(\mathcal{I},\mathcal{D}^{b}_{\mathrm{dg}}\left(\mathcal{C}\right)\right)\left(\iota,\psi_{M}\circ\iota\right)\cong\mathrm{Ext}^{*}_{\mathcal{C}_{\mathcal{I}}}\left(\iota,\psi_{M}\circ\iota\right).$$c_{\iota_{I},I}=\iota_{I,*}$
The above observation is in particular useful as explicit computations for bimodules between $𝕜\Bbbk$-linear categories, or, even more, $\mathcal{A}_{\infty}$/$\mathrm{dg}$-categories can be very unwieldy. However, the above observation allows us to compute the equivariant characteristic morphisms of the $I$-$\mathcal{C}$ bimodule $\iota_{I}$ in special cases.
3.2. The $\mathrm{Inj}$-equivariant characteristic morphism
We use the observations on the structure of the characteristic morphism from the last section in order to compute the equivariant characteristic morphism of the inclusion of injective objects. In particular we consider throughout this section the functor
$\iota:\mathrm{Inj}\hookrightarrow\mathcal{C}\subset\mathcal{D}_{\mathrm{dg}}\left(\mathcal{C}\right)$
Lemma 3.3.
Let $\mathcal{C}$ be a $𝕜\Bbbk$-linear category with enough injectives. Then the inclusion $\iota:\mathrm{Inj}\hookrightarrow\mathcal{C}\subset\mathcal{D}_{\mathrm{dg}}\left(\mathcal{C}\right)$ is dg.
Proof.
Let $I,J\in\mathrm{Inj}$. Then $\iota\left(I\right)$ and $\iota\left(J\right)$ are injective. In particular we have
$\mathcal{D}^{b}_{\mathrm{dg}}\left(\mathcal{C}\right)\left(\iota I,\iota J\right)$ $=\mathcal{D}^{b}_{\mathrm{dg}}\left(\mathcal{C}\right)\left(I,J\right)$
$=\mathcal{C}\left(I,J\right)$
as graded vector-spaces concentrated in degree 00. In particular
$\iota:\mathrm{Inj}\hookrightarrow\mathcal{D}^{b}_{\mathrm{dg}}\left(\mathrm{Qcoh}\left(X\right)\right)$
is dg. ∎
Proposition 3.4.
Let $\mathcal{C}$ be a $𝕜\Bbbk$-linear category with enough injectives. Then the functor $\iota^{*}:\mathcal{D}_{\mathrm{dg}}\left(\mathcal{C}\right)\to\mathcal{D}_{\mathrm{dg}}\left(\mathrm{Inj}\right)$ is a quasi-equivalence.
Proof.
Since $\mathcal{C}$ has finite global dimension every object in $\mathcal{D}^{b}\left(\mathcal{C}\right)$ has a finite resolution of injectives. In particular every object $N\in\mathcal{D}^{b}_{\mathrm{dg}}\left(\mathcal{C}\right)$ is quasi-isomorphic to a bounded complex of injectives. These complexes are dg-$\mathrm{Inj}$-modules and so $\iota:\mathrm{Inj}\hookrightarrow\mathcal{C}$ induces an equivalence
$\iota_{*}:\mathcal{D}^{b}_{\mathrm{dg}}\left(\mathcal{C}\right)\to\mathcal{D}^{b}_{\mathrm{dg}}\left(\mathrm{Inj}\right)$
as claimed. ∎
Corollary 3.5.
Let $\mathcal{C}$ be a $𝕜\Bbbk$-linear category with enough injectives and finite global dimension. Then the equivariant characteristic morphism $c_{\iota,\mathrm{Inj}}$ is an isomorphism.
Proof.
By Section 3.1 above we can interpret the equivariant characteristic morphism as the pullback along the inclusion of the full subcategory $\iota:\mathrm{Inj}\to\mathcal{D}^{b}_{\mathrm{dg}}\left(\mathcal{C}\right)$ in the homotopy category of $\mathrm{dg}$-categories. In particular we need to pick a fibrant replacement of $\mathrm{Inj}$. By [Tab07, Lemme 2.13] pretriangulated replacements are an incarnation of a fibrant replacement. So we may choose the canonical pretriangulated hull given by $\mathrm{dg}$-modules $\mathcal{D}_{\mathrm{dg}}\left(\mathrm{Inj}\right)$. By Proposition 3.4 $\iota$ induces an equivalence
$\iota_{*}:\mathcal{D}_{\mathrm{dg}}\left(\mathcal{C}\right)\to\mathcal{D}_{\mathrm{dg}}\left(\mathrm{Inj}\right)$
and hence in particular $c_{\iota,\mathrm{Inj}}\cong\iota_{*}$ is an isomorphism. ∎
Corollary 3.6.
Let $X$ be a smooth finite dimensional scheme and consider the category $\mathrm{Qcoh}\left(X\right)$. Then the equivariant characteristic morphism $c_{\iota,\mathrm{Inj}}$ is an isomorphism.
Proof.
We have that $\mathrm{Qcoh}\left(X\right)$ has enough injectives. In particular by Corollary 3.5 we have that
$c_{\iota,\mathrm{Inj}}:\mathrm{HH}^{*}\left(\mathrm{Qcoh}\left(X\right),\mathcal{M}\right)\xrightarrow{\sim}\mathrm{Ho}\left(\mathrm{dg}-\mathrm{Cat}\right)\left(\mathrm{Inj},\mathcal{D}^{b}_{\mathrm{dg}}\left(\mathrm{Qcoh}\left(X\right)\right)\right)\left(\iota,\psi_{\mathcal{M}}\circ\iota\right)$
is an isomorphism. ∎
Remark 3.7.
The above discussion shows that for smooth projective schemes the full subcategory of injectives behaves like a tilting object.
Corollary 3.8.
Let $\mathcal{C}$ be a $𝕜\Bbbk$-linear category with enough injectives, $\mathcal{M}$ a $𝕜\Bbbk$-central $\mathcal{X}$-bimodule and let $0\neq\eta\in\mathrm{HH}^{*}\left(\mathcal{C},\mathcal{M}\right)$. Then we have that its image under the $\mathrm{Inj}$-equivariant characteristic morphism $c_{\iota,\mathrm{Inj}}\left(\eta\right)$ does not vanish.
4. Construction of $\psi_{\eta}$
We recall the construction of the candidate non-Fourier-Mukai functor $\psi_{\eta}$. In particular from now on we will exclusively consider a functor of the shape $\psi_{\eta}$ for a smooth divisor $X\hookrightarrow Y$ with $\dim X=n$ and $0\neq\eta\in\mathrm{HH}^{m}\left(X,M\right)$ with $M$ a line-bundle, $m\geq n+3$ and $f_{*}\eta=0$.
In order to apply Definition 4.1, Lemma 4.2 and Lemma 4.4 we assume that every quasi-projective scheme $X$ comes equipped with an open affine cover
$X=\bigcup_{i=0}^{k}U_{i}.$
We recall the following construction originally introduced by W. Lowen and M. Van den Bergh [LVdB11].
Definition 4.1.
[LVdB11, Definition 4.2] Let $X=\bigcup_{i=0}^{k}U_{i}$ be an open affine cover of a quasiprojective scheme $X$. Consider for $I\subset\left\{1,...,k\right\}$ the open sets $U_{I}:=\bigcap_{i\in I}U_{i}$. Then $\mathcal{X}$ is the category with objects $\mathrm{P}\left\{1,...,k\right\}\setminus\emptyset$, morphisms:
$\mathcal{X}\left(I,J\right):=\begin{cases}\mathcal{O}_{X}\left(U_{J}\right)&I\subset J\\ 0&\text{else}\end{cases}$
and composition of morphisms induced by multiplication in the local ring composed with the restriction morphism.
One can think about $\mathcal{X}\mathrm{-mod}$ roughly as the category of presheaves associated to an affine open cover. It comes with the following useful property which is a collection of results from [RVdBN19] and which originally appeared in [LVdB11].
Lemma 4.2.
[Kü22, Lemma 5.2] Let $X$ be smooth quasi-projective. Then there is a fully faithful embedding
$w:\mathcal{D}^{b}\left(\mathrm{Qcoh}X\right)\hookrightarrow\mathcal{D}^{b}\left(\mathcal{X}\right)$
with essential image given by $\mathcal{D}^{b}_{w\mathrm{Qcoh}X}\left(\mathcal{X}\right)$ and a fully faithful embedding
$W:\Delta_{*}\mathcal{D}\left(\mathrm{Qcoh}X\right)\hookrightarrow\mathcal{D}\left(\mathcal{X}\otimes\mathcal{X}^{\mathrm{op}}\right),$
where $\Delta_{*}\mathcal{D}\left(\mathrm{Qcoh}X\right)$ is the essential image of the direct image of the diagonal $\Delta:X\hookrightarrow X\times X$. In particular we have for quasi-coherent $M$
$\mathrm{HH}^{*}\left(X,M\right)\cong\mathrm{HH}^{*}\left(\mathcal{X},WM\right).$
Remark 4.3.
We will denote $WM$ by $\mathcal{M}$ for readability, similar to $\mathcal{X}$ for the $𝕜\Bbbk$-linear category associated to $X$. Observe in particular that we have $W\mathcal{O}_{\Delta}\cong\mathcal{X}$ giving rise to the last equality in Lemma 4.2.
Lemma 4.4.
[Kü22, Lemma 5.4] Let $f:X\hookrightarrow Y$ be a closed embedding of quasi projective schemes. Then we may pull back the open affine cover of $Y$ in the construction of $\mathcal{Y}$ and get an induced functor $\mathfrak{f}:\mathcal{Y}\to\mathcal{X}$ such that the diagram
$\mathcal{D}^{b}\left(\mathcal{X}\right)$$\mathcal{D}^{b}\left(\mathrm{Qcoh}X\right)$$\mathcal{D}^{b}\left(\mathrm{Qcoh}Y\right)$$\mathcal{D}^{b}\left(\mathcal{Y}\right),$$\mathfrak{f}_{*}$$w$$w$${f}_{*}$
commutes.
As we will mostly be working with the category $\mathcal{D}^{b}\left(\mathcal{X}\right)$ and $w$ is an exact embedding we will denote the functor $\mathfrak{f}:\mathcal{X}\to\mathcal{Y}$ by $f:\mathcal{X}\to\mathcal{Y}$ for ease of readability. We also use the slightly strange convention of $f_{*}$ for the pullback on the $𝕜\Bbbk$-linear side in order to mimic the geometric setting and improve readability.
The following is an application of [RVdBN19, Proposition 5.3.1] to our setting.
Proposition 4.5.
[Kü22, Proposition 5.5] Let $X$ be smooth projective of dimension $n$ and let $\eta\in\mathrm{HH}^{\geq n+3}\left(X,M\right)$. Then there exists an exact functor
$L:\mathcal{D}^{b}\left(\mathrm{Qcoh}X\right)\to\mathcal{D}^{b}_{wQcoh\left(X\right)}\left(\mathcal{X}_{\eta}\right).$
We will study the functor arising from the following Lemma 4.6 in more detail in Section 6. However, since we want to recall here the basic ingredients to the construction we include its basic construction.
Proposition 4.6.
[RVdBN19, Proposition 8.2.6] Let $f:\mathcal{Y}\to\mathcal{X}$ be a $𝕜\Bbbk$-linear functor between $𝕜\Bbbk$-linear categories $\mathcal{X}$ and $\mathcal{Y}$, $\mathcal{M}$ a $𝕜\Bbbk$-central $\mathcal{X}$-bimodule and $\eta\in\mathrm{HH}^{\geq 3}\left(\mathcal{X},\mathcal{M}\right)$ such that $0=f_{*}\eta\in\mathrm{HH}^{*}\left(\mathcal{Y},f_{*}\mathcal{M}\right)$. Then there exists an $\mathcal{A}_{\infty}$-functor $\widetilde{f}$ making the following diagram commute
[…]
Construction 4.7.
[RVdBN19] Let $X\hookrightarrow Y$ be the inclusion of a smooth divisor into a smooth projective scheme and let
$0\neq\eta\in\ker\left(f_{*}:\mathrm{HH}^{m}\left(X,M\right)\to\mathrm{HH}^{m}\left(Y,f_{*}M\right)\right)$
with $m\geq\dim X+3$.
Then we can construct a functor
[…]
where the functor $L$ is given by Proposition 4.5 and $\widetilde{f}_{*}$ exists by Proposition 4.6.
Remark 4.8.
Construction 4.6 works out as well if $\eta$ is zero, however, the arising functor then trivially admits a lift to the $\mathcal{A}_{\infty}$-level and so we will not consider this case.
5. Recollection on Obstructions
As we will obstruct $\mathcal{A}_{\infty}$- respectively $\mathrm{dg}$-lifts we recall the following notions and results from [RVdBN19] on colifts along infinitesimal deformations and $\mathcal{A}_{\infty}$-obstructions:
Definition 5.1.
Let $\mathcal{X}$ be a small $𝕜\Bbbk$-linear category, assume $\mathcal{M}$ is right-flat, $\eta\in\mathrm{HH}^{\geq 3}\left(\mathcal{X},\mathcal{M}\right)$ and let $U\in\mathcal{X}\mathrm{-mod}$.
A lift of $U$ to $\mathcal{X}_{\eta}$ is an object $V\in\mathcal{D}^{b}_{\mathrm{dg}}\left(\mathcal{X}_{\eta}\right)$, such that there exists an isomorphism in $\mathcal{X}_{\eta}\mathrm{-mod}$:
$\phi:\mathrm{H}^{*}\left(V\right)\xrightarrow{\sim}\mathrm{H}^{*}\left(\mathcal{X}_{\eta}\right)\otimes_{\mathcal{X}}U.$
Dually, if $\mathcal{M}$ is left-projective and $U\in\mathcal{X}\mathrm{-mod}$. A colift of $U$ to $\mathcal{X}_{\eta}$ is an object $V\in\mathcal{D}^{b}_{\mathrm{dg}}\left(\mathcal{X}_{\eta}\right)$, such that there exists an isomorphism in $\mathcal{X}_{\eta}\mathrm{-mod}$:
$\phi:\mathrm{H}^{*}\left(V\right)\xrightarrow{\sim}\mathrm{Hom}_{\mathcal{X}}\left(\mathrm{H}^{*}\left(\mathcal{X}_{\eta}\right),U\right).$
Theorem 5.2 ([RVdBN19, Lemma 8.2.1]).
The object $U\in\mathcal{X}\mathrm{-mod}$ has a lift to $\mathcal{X}_{\eta}$ if and only if $c_{U}\left(\eta\right)=0$. It has a colift if and only if $c^{*}_{U}\left(\eta\right)=0$.
For the next Lemma 5.3 observe that if $M$ is a line-bundle we have that the bimodule $\mathcal{M}:=WM$ is an invertible $\mathcal{X}$-bimodule.
Lemma 5.3 ([RVdBN19, Lemma 6.3.1]).
Assume $\mathcal{M}$ is an invertible $\mathcal{X}$-bimodule, then we have that $c_{U}\left(\eta\right)\neq 0$ if and only if $c^{*}_{U}\left(\eta\right)\neq 0$.
Lemma 5.4 ([RVdBN19, Lemma 8.2.1]).
Let $f_{i}:\mathcal{A}\to\mathcal{B}$ be an $\mathcal{A}_{i}$-functor between two $\mathcal{A}_{\infty}$-categories. Then there are inductive obstructions
$o_{i+1}\left(f_{i}\right)\in\mathrm{HH}^{i+1}\left(\mathrm{H}^{*}\left(\mathcal{A}\right),\mathrm{H}^{*}\left(\mathcal{B}\right)\right)_{1-i}$
such that $o_{i+1}\left(f_{i}\right)$ vanishes if and only if there exists a $\delta$ with $[\delta,m_{1}]=0$ such that $f_{i}+\delta$ can be lifted to an $\mathcal{A}_{i+1}$-morphism.
This obstruction is natural i.e. if we have two $\mathcal{A}_{\infty}$-functors $G:\mathcal{A}^{\prime}\to\mathcal{A}$ and $G^{\prime}:\mathcal{B}\to\mathcal{B}^{\prime}$ we get
$o_{i+1}\left(G^{\prime}\circ f_{i}\circ G\right)=\mathrm{H}^{*}\left(G^{\prime}\right)\circ o_{i+1}\left(f_{i}\right)\circ\mathrm{H}^{*}\left(G\right).$
The obstructions in Lemma 5.4 are given by the $\mathcal{A}_{\infty}$-functor formulas and the observation that they naturally define Hochschild cohomology classes. In particular they are given by the formulas
(2) $o_{i}\left(f\right)=\sum_{k=1}^{i}\pm\left[\mathrm{m}_{k},f_{i-k+1}\right]\in\mathrm{HH}^{i+1}\left(\mathrm{H}^{*}\left(\mathcal{A}\right),\mathrm{H}^{*}\left(\mathcal{B}\right)\right)_{1-i}$
Lemma 5.5.
Let $\iota\in\mathcal{D}\left(\mathcal{X}\otimes\mathcal{I}^{op}\right)$ for a small $𝕜\Bbbk$-linear category $\mathcal{I}$. Let $\widehat{\iota}\in\mathcal{D}_{\mathrm{dg}}\left(\mathcal{X}_{\eta}\otimes\mathcal{I}^{op}\right)$. Then the following are equivalent:
  • The object $\widehat{\iota}$ is a colift of $\iota\in\mathcal{D}\left(\mathcal{X}\otimes\mathcal{I}^{op}\right)$.
  • The object $\widehat{\iota}$ is an $\mathcal{A}_{\infty}$-lift of $L\circ\iota\in\mathcal{D}\left(\mathcal{X}_{\eta}\right)_{\mathcal{I}}$ i.e. $L\circ\iota:\mathcal{I}\to\mathcal{X}_{\eta}$.
Proof.
The object $\widehat{\iota}$ is a colift of $\iota$ if and only if we have an isomorphism of $\mathrm{H}^{*}\left(\mathcal{X}_{\eta}\right)$-modules:
$\Phi:\mathrm{Ext}^{*}_{\mathcal{X}\otimes\mathcal{I}^{op}}\left(\left(\mathcal{X}\otimes\mathcal{I}^{op}\right)_{\eta\cup 1},\iota\right)\xrightarrow{\sim}\mathrm{H}^{*}\left(\widehat{\iota}\right).$
In particular it suffices, since by Remark 2.7 $\mathcal{X}_{\eta}\otimes\mathcal{I}^{op}\cong\left(\mathcal{X}\otimes\mathcal{I}^{op}\right)_{\eta\cup 1}$, to find an isomorphism
$\mathrm{Ext}^{*}_{\mathcal{X}\otimes\mathcal{I}^{op}}\left(\left(\mathcal{X}\otimes\mathcal{I}^{op}\right)_{\eta\cup 1},\iota\right)\xrightarrow{\sim}L\circ{\iota},$
as then we can identify the two properties.
For this consider the following concatenation of natural isomorphisms:
$\mathrm{Ext}^{*}_{\mathcal{X}\otimes\mathcal{I}^{op}}\left(\left(\mathcal{X}\otimes\mathcal{I}^{op}\right)_{\eta\cup 1},\iota\right)$ $\cong\mathrm{Ext}^{*}_{\mathcal{X}\otimes\mathcal{I}^{op}}\left(\mathcal{X}_{\eta}\otimes\mathcal{I}^{op},\iota\right)$
$\cong\mathrm{Ext}^{*}_{\mathcal{X}}\left(\mathcal{X}_{\eta},\iota\right)$
$\cong\mathrm{Ext}^{*}_{\mathcal{X}_{\eta}}\left(\mathcal{X}_{\eta},L\circ\iota\right)$ [RVdBN19, Corollary 5.3.2]
$\cong L\circ\iota.$
So we have that if $\widehat{\iota}$ is a colift, it immediately defines by the above isomorphisms an $\mathcal{A}_{\infty}$-lift of $L\circ\iota$ and vice versa. ∎
Proposition 5.6.
Let $\iota:\mathrm{Inj}\hookrightarrow\mathcal{D}^{b}\left(\mathrm{Qcoh}X\right)$ be the inclusion of injectives in the canonical heart. Then the functor $L\circ\iota:\mathrm{Inj}\to\mathcal{D}^{b}\left(\mathcal{X}_{\eta}\right)$ does not admit an $\mathcal{A}_{\infty}$-lift.
Proof.
By Lemma 5.5 we have that a lift of $L\circ\iota$ would be a colift of the $\mathrm{Inj}$-$\mathcal{X}$ bimodule $\iota$. This colift is obstructed by the dual characteristic morphism $c^{*}_{\iota,\mathrm{Inj}}\left(\eta\right)$. As $\mathcal{M}$ is invertible, we have by Lemma 5.3 that the dual characteristic morphism vanishes if and only if the characteristic morphism vanishes. By Corollary 3.8 we have that $c_{\iota,\mathrm{Inj}}\left(\eta\right)\neq 0.$
In particular such a colift cannot exist, and hence no $\mathcal{A}_{\infty}$-lift as well. ∎
6. Structure of the functor $\widetilde{f}$
It will be key for our conclusion that we carefully pick the space we consider for computing our obstructions, respectively understand the structure of the functor $\widetilde{f}$. In particular we will from now on fix the inclusion of a smooth divisor $f:X\hookrightarrow Y$, a pulled back linebundle $M=f^{*}\widehat{M}$ and $0\neq\eta\in\mathrm{HH}^{m}\left(X,M\right)$ with $m\geq\dim X+3$.
We start of by computing the images of sheaves $\mathcal{F}\in\mathrm{Qcoh}\left(X\right)$ under the functor $\psi_{\eta}$:
Lemma 6.1.
Let $F\in\mathrm{Qcoh}\left(X\right)$. Then
$\psi_{\eta}F\cong f_{*}F\oplus\left(f_{*}F\otimes f_{*}M\left[2-m\right]\right).$
Proof.
Consider the distinguished triangle in $\mathcal{D}^{b}\left(\mathcal{X}_{\eta}\right)$ from [RVdBN19, Lemma 11.3]
$\mathrm{can}_{*}\mathcal{F}\to L\mathcal{F}\to\mathrm{can}_{*}\Sigma^{2-m}\left(\mathcal{F}\otimes\mathcal{M}\right)\to\Sigma\mathrm{can}_{*}\mathcal{F},$
where we interpret $\mathcal{F}$ as an object in $\mathcal{D}^{b}\left(\mathcal{X}_{\eta}\right)$ via the functor $\mathrm{can}_{*}$. Applying the exact functor $\widetilde{f}_{*}$ from Lemma 4.6 we get the following triangle in $\mathcal{D}^{b}\left(\mathcal{Y}\right)$
$\left(\widetilde{f}\circ\mathrm{can}\right)_{*}\mathcal{F}\to\psi_{\eta}\mathcal{F}\to\left(\widetilde{f}\circ\mathrm{can}\right)_{*}\Sigma^{2-m}\left(\mathcal{F}\otimes\mathcal{M}\right)\to\left(\widetilde{f}\circ\mathrm{\mathrm{can}}\right)_{*}\Sigma\mathcal{F}.$
As $\mathrm{can}\circ\widetilde{f}\cong f$ the triangle turns into
$f_{*}\mathcal{F}\to\psi_{\eta}\mathcal{F}\to f_{*}\Sigma^{2-m}\left(\mathcal{F}\otimes\mathcal{M}\right)\to f_{*}\Sigma\mathcal{F}.$
And since $f_{*}$ is exact we have that it is in particular graded. So we have the triangle
$f_{*}\mathcal{F}\to\psi_{\eta}\mathcal{F}\to\Sigma^{2-m}f_{*}\left(\mathcal{F}\otimes\mathcal{M}\right)\to\Sigma f_{*}\mathcal{F}.$
Now, as $\mathcal{F}$ and $\mathcal{M}$ are sheaves we have
$\mathrm{Ext}^{1}_{\mathcal{Y}}\left(\Sigma^{2-m}f_{*}\mathcal{F}\otimes\mathcal{M},f_{*}\mathcal{F}\right)$ $\cong\mathrm{Ext}^{m-1}_{\mathcal{Y}}\left(f_{*}\mathcal{F}\otimes\mathcal{M},f_{*}\mathcal{F}\right)$ $\mathrm{Ext}^{i}\left(\Sigma^{-j}\_,\_\right)\cong\mathrm{Ext}^{i+j}\left(\_,\_\right)$
$\cong\mathrm{Ext}^{m-1}_{Y}\left(f_{*}\mathcal{F}\otimes\mathcal{M},f_{*}\mathcal{F}\right)$ Lemma 4.2
[…]
where we used in the last line that $m\geq\dim X+3$ gives $m-1\geq\dim Y=\dim X+1$. So the triangle
$f_{*}\mathcal{F}\to\psi_{\eta}\mathcal{F}\to f_{*}\Sigma^{2-m}\mathcal{F}\otimes\mathcal{M}\to\Sigma\mathcal{F}$
splits and we have
$\psi_{\eta}\left(\mathcal{F}\right)\cong f_{*}\mathcal{F}\oplus f_{*}\left(\Sigma^{2-m}M\otimes\mathcal{F}\right),$
as claimed. ∎
Corollary 6.2.
We have that the space $\mathrm{Ext}^{*}\left(\psi_{\eta}I,\psi_{\eta}J\right)$ is concentrated in degrees $m-2,m-3,1,0,3-m$ and $2-m$ for $I,J\in\mathrm{Inj}$.
Proof.
By Lemma 6.1 we have $\psi_{\eta}I\cong f_{*}I\oplus f_{*}\left(I\otimes M\left[2-m\right]\right)$ which by projection formula and $M\cong f^{*}\widehat{M}$ turns into $f_{*}I\oplus f_{*}\left(I\otimes\widehat{M}\right)\left[2-m\right]$. In particular we can compute
$\mathrm{Ext}_{Y}^{*}\left(\psi_{\eta}I,\psi_{\eta}J\right)$ $\cong\mathrm{Ext}_{Y}^{*}\left(f_{*}I\oplus\left(f_{*}I\otimes f_{*}M\left[2-m\right]\right),f_{*}J\oplus f_{*}\left(J\otimes M\left[2-m\right]\right)\right)$
$\cong\begin{pmatrix}\mathrm{Ext}_{Y}^{*}\left(f_{*}I,f_{*}J\right)&\mathrm{Ext}_{Y}^{*}\left(f_{*}I,f_{*}\left(J\otimes M\right)[2-m]\right)\\ \mathrm{Ext}_{Y}^{*}\left(f_{*}\left(I\otimes M\right)[2-m],f_{*}J\right)&\mathrm{Ext}_{Y}^{*}\left(f_{*}I,f_{*}J\right)\end{pmatrix}$
Where we applied in the bottom right corner the fact that $\widehat{M}$ is invertible while also dropping the shifts. Now it suffices to observe that we have
$Ext_{Y}^{*}\left(f_{*}F,f_{*}G\right)$ $\cong\mathrm{Ext}^{*}_{X}\left(f^{*}f_{*}F,G\right)$
$\cong\mathrm{Ext}^{*}_{X}\left(I\otimes C_{f}[-1]\oplus F,G\right)$ $f^{*}f_{*}F\cong\left(F\oplus F\otimes C_{f}[1]\right)$
$\cong\mathrm{Ext}^{*}_{X}\left(F,G\right)\oplus\mathrm{Ext}^{*}_{X}\left(F\otimes C_{f},G\right)$ $\mathrm{Ext}^{*}\text{ is additive}$
Where we use $C_{f}$ for the conormal bundle of the inclusion $f:X\hookrightarrow Y$. Now it suffices to observe that as $f$ is a divisor $C_{f}$ is invertible too and in our cases $F$ and $G$ are either injective, or an injective tensored by an invertible module $M$. ∎
In order to have control over the obstructions explicitly, we will construct the functor $\widetilde{f}:\mathcal{Y}\to\mathcal{X}_{\eta}$ from Lemma 4.6 explicitly. For this consider the commutative diagram of $\mathcal{A}_{\infty}$-categories and functors giving rise to a good candidate.
[…]
The bottom arrow is the functor given in Lemma 4.4 and we denoted by can the canonical functors arising from the infinitesimal deformation.
The other functors are given by:
  • The functor $f:\mathcal{Y}_{f_{*}\eta}\to\mathcal{X}_{\eta}$ is the functor given by the identity on objects and acting on morphisms by
    $\begin{pmatrix}f&0\\ 0&f\end{pmatrix}:\mathcal{Y}_{f_{*}\eta}\left(i,j\right)=\mathcal{Y}\left(i,j\right)\oplus f_{*}\mathcal{M}\left(i,j\right)[2-m]\to\mathcal{X}\left(i,j\right)\oplus\mathcal{M}\left(i,j\right)[2-m].$
    This functor is by construction an $\mathcal{A}_{\infty}$-functor as by definition $f_{*}\eta$ is given by the composition of bimodule morphisms
[…]
  • The functor $h:\mathcal{Y}_{0}\to\mathcal{Y}_{f_{*}\eta}$ is given by a nullhomotopy of
    $0=f_{*}\eta\in\mathrm{HH}^{*}\left(\mathcal{Y},f_{*}M\right)$
    from Proposition 2.6 and is in particular an $\mathcal{A}_{\infty}$-equivalence.
  • The functor $\mathcal{Y}\oplus\mathcal{M}[2-m]\to\mathcal{Y}_{0}$ is the definition of the trivial infinitesimal deformation.
The diagram (3) now shows that the main choice involved in the definition of $\widetilde{f}$ is the choice of a split $s:\mathcal{Y}\to\mathcal{Y}\oplus f_{*}\mathcal{M}[2-m]$. For which we take the one given by
$s=\begin{pmatrix}\mathrm{Id}\\ 0\end{pmatrix}:\mathcal{Y}\to\mathcal{Y}\oplus f_{*}\mathcal{M}[2-m].$
In particular we can use these constructions to study the functors involved in the construction of $\widetilde{f}$ explicitly and conclude that we can restrict to a suitable space where we can control the $\mathcal{A}_{\infty}$-bimodule structure rigorously.
Definition 6.3.
We consider for the remainder of this paper the concrete choice of the functor $\widetilde{f}:\mathcal{Y}\to\mathcal{X}_{\eta}$ given by the composition
[…]
Now we continue our study by restricting to the category of injectives in $\mathcal{D}^{b}\left(\mathrm{Qcoh}X\right)$,
$\iota:\mathrm{Inj}\hookrightarrow\mathcal{D}^{b}\left(\mathrm{Qcoh}X\right).$
For this consider the following composition of all intermediate steps of the construction of $\psi_{\eta}$. We also include a step for the essential image of $\mathrm{Inj}$ in $\mathcal{D}^{b}\left(\mathcal{X}_{\eta}\right)$ respectively in $\mathcal{D}^{b}\left(\mathcal{Y}\right)$, which we denote by $\mathrm{Inj}_{\eta}$, respectively $f_{*}\mathrm{Inj}_{f_{*}\eta}$,
[…]
Remark 6.4.
[…]
We then use the explicit structure of the category $f_{*}\mathrm{Inj}_{f_{*}\eta}$ to compute that these obstructions cannot vanish independent of intermediate lifts.
Conceptually one can think about this argument as the bimodule $\psi_{\eta}$ not being $\mathcal{A}_{\infty}$ over $\mathrm{Inj}$, however, it is $\mathcal{A}_{\infty}$ over $\mathcal{Y}$. In particular it behaves like a non-$\mathcal{A}_{\infty}$ Fourier-Mukai kernel
The following Propositions 6.5 and Proposition 6.6 shows that the notations $\mathrm{Inj}_{\eta}$ and $f_{*}\mathrm{Inj}_{f_{*}\eta}$ for the essential image of $f_{*}\circ L\circ\iota$ are warranted.
Proposition 6.5.
The essential image of $L\circ\iota$ has objects $I\in\mathrm{Inj}$, morphism spaces given by
$\textrm{essim}\left(L\circ\iota\right)\left(I,J\right)\cong\mathrm{Ext}_{X}^{0}\left(I,J\right)\oplus\mathrm{Ext}_{X}^{0}\left(I,J\otimes\mathcal{M}\right)[2-m]$
and $\mathrm{m}_{m}=\eta$.
Proof.
By construction of $L$, in particular [RVdBN19, (5.7)], the morphism spaces in $\mathrm{Inj}_{\eta}$ are of the shape
$\mathrm{essim}\left(L\circ\iota\right)\left(I,J\right)$ $\cong\mathrm{Ext}_{X}^{*}\left(I,J\right)\oplus\mathrm{Ext}_{X}^{2-m+*}\left(I,J\otimes\mathcal{M}\right)$
$\cong\mathrm{Ext}_{X}^{0}\left(I,J\right)\oplus\mathrm{Ext}_{Y}^{0}\left(I,J\otimes\mathcal{M}\right)[2-m].$
As $L$ is defined by identifying the inclusion of $\mathrm{Inj}$ into $\mathcal{D}^{b}\left(\mathcal{X}\right)$ with the inclusion into $\mathcal{D}^{b}\left(\mathcal{X}_{\eta}\right)$ [RVdBN19, Lemma 6.3.1] we have that there is a non-trivial $\mathrm{m}_{m}=\eta$, via the identification from [LVdB05, Theorem 8.17] $\mathrm{HH}^{*}\left(\mathrm{Qcoh},\mathcal{M}\right)\cong\mathrm{HH}^{*}\left(\mathrm{Inj},\mathcal{M}\right)$. ∎
Proposition 6.6.
The essential image of $f_{*}\circ L\circ\iota$ has objects $i\in I$ and morphism spaces given by
$\textrm{essim}\left(f_{*}\circ L\circ\iota\right)\left(I,J\right)\cong\mathrm{Ext}_{Y}^{*}\left(f_{*}I,f_{*}J\right)\oplus\mathrm{Ext}_{Y}^{2-m+*}\left(f_{*}I,f_{*}\left(J\otimes\mathcal{M}\right)\right).$
Proof.
As $f_{*}$ is an $\mathcal{A}_{\infty}$-functor it suffices to push forward the structure from Proposition 6.5. ∎
Proposition 6.7.
The functor $s_{*}:\mathcal{D}^{b}\left(Y_{0}\right)\to\mathcal{D}^{b}\left(\mathcal{Y}\right)$ acts on objects of the shape $h_{*}\circ f_{*}\circ L\left(I\right)$ for $I\in Inj$ by
$s_{*}:\begin{pmatrix}\mathrm{Ext}_{Y}^{*}\left(f_{*}I,f_{*}J\right)\\ \mathrm{Ext}_{Y}^{*+m-2}\left(f_{*}I,f_{*}\left(J\otimes M\right)\right)\end{pmatrix}$ $\to\mathrm{Ext}_{Y}^{*}\left(f_{*}I\oplus\Sigma^{m-2}f_{*}\left(I\otimes M\right),f_{*}J\oplus\Sigma^{m-2}f_{*}\left(J\otimes M\right)\right)$
$\begin{pmatrix}\alpha\\ \beta\end{pmatrix}$ $\mapsto\begin{pmatrix}\alpha&\beta\\ 0&\alpha\end{pmatrix}$
Proof.
By the above discussions and Proposition 6.6 we have
$\mathrm{essim}\left(f_{*}\circ L\circ\iota\right)\left(I,J\right)\cong\mathrm{Ext}_{Y}^{*}\left(f_{*}I,f_{*}J\right)\oplus\mathrm{Ext}_{Y}^{2-m+*}\left(f_{*}I,f_{*}\left(J\otimes\mathcal{M}\right)\right).$
As $h$ is a homotopy equivalence the same holds for $\mathrm{essim}\left(h_{*}\circ f_{*}\circ L\circ\iota\right)$. By Lemma 6.1 we have $\psi_{\eta}\left(I\right)\cong f_{*}I\oplus f_{*}\left(M\otimes I\right)[2-m]$. As we made the choice for $s$ being the trivial inclusion $\mathcal{Y}\hookrightarrow\mathcal{Y}_{0}\cong\mathcal{Y}\oplus f_{*}\mathcal{M}[2-m]$, $s_{*}$ induces the inclusions of upper triangular matrices with identical diagonal, as claimed. ∎
Corollary 6.8.
The functor $s_{*}:\mathcal{D}^{b}\left(\mathcal{Y}_{0}\right)\to\mathcal{D}^{b}\left(\mathcal{Y}\right)$ is an isomorphism on negative degree morphisms in the essential image of $\mathrm{Inj}$.
Proof.
By Proposition 6.7 we have that in the matrix expression of
$\begin{pmatrix}\alpha&\beta\\ \gamma&\delta\end{pmatrix}\in\mathrm{Ext}_{Y}^{*}\left(f_{*}I\oplus\Sigma^{m-2}f_{*}\left(I\otimes M\right),f_{*}J\oplus\Sigma^{m-2}f_{*}\left(J\otimes M\right)\right)$
the morphisms $\alpha,\delta$ and $\gamma$ must be in non-negative degrees as the direct summands $\mathrm{Ext}^{*}_{Y}\left(f_{*}I,f_{*}J\right)$,$\mathrm{Ext}_{Y}^{*}\left(f_{*}I\otimes M,f_{*}J\otimes M\right)$ and $\mathrm{Ext}_{Y}^{*}\left(\Sigma^{m-2}f_{*}I\otimes M,f_{*}J\right)$ are concentrated in non-negative degrees.
In particular it suffices to verify the claim for $\beta$. There we have that the map
$s_{*}:\mathrm{Ext}_{Y}^{*+m-2}\left(f_{*}I,f_{*}\left(J\otimes M\right)\right)$ $\to\mathrm{Ext}_{Y}^{*}\left(f_{*}I\oplus\Sigma^{m-2}f_{*}\left(I\otimes M\right),f_{*}I\oplus\Sigma^{m-2}f_{*}\left(I\otimes M\right)\right)$
$\beta$ $\mapsto\begin{pmatrix}0&\beta\\ 0&0\end{pmatrix}$
is an isomorphism given by the inclusion of the direct summand
$\mathrm{Ext}_{Y}^{*}\left(f_{*}I,\Sigma^{m-2}f_{*}\left(J\otimes M\right)\right)\hookrightarrow\mathrm{Ext}_{Y}^{*}\left(f_{*}I\oplus\Sigma^{m-2}f_{*}\left(I\otimes M\right),f_{*}I\oplus\Sigma^{m-2}f_{*}\left(I\otimes M\right)\right).$
After restricting to the negative morphisms this remains an isomorphism, which finishes the claim. ∎
7. Computing obstructions explicitly for Divisors
The observations from the previous section let us compute:
Proposition 7.1.
The obstruction
$o_{m}\left(L\circ\iota\right)=[m_{m},L]\in\mathrm{HH}^{*}\left(\mathrm{Inj},\mathrm{Inj}_{\eta}\right)_{2-m}$
does not vanish.
Proof.
We know by Proposition 5.6 that $L\circ\iota$ does not have an $\mathcal{A}_{\infty}$-lift. In particular we have by Lemma 5.4 that there must be a non-vanishing
$o_{i}\left(L\circ\iota\right)\in\mathrm{HH}^{i}\left(\mathrm{Inj},\mathrm{Inj}_{\eta}\right)_{i}$
However, as $L\circ\iota$ is a functor it is $\mathcal{A}_{2}$ and the non-vanishing obstruction must have $i\geq 3$. As we have by Proposition 6.5 that $\mathrm{Inj}_{\eta}$ is concentrated in degrees 00 and $2-m$ we have that the only non-trivial obstruction can be
$o_{m}\left(L\circ\iota\right)\in\mathrm{HH}^{*}\left(\mathrm{Inj},\mathrm{Inj}_{\eta}\right)_{2-m}$
and in particular $o_{m}\left(L\circ\iota\right)\neq 0$. ∎
Lemma 7.2.
The obstructions of $\mathcal{A}_{i}$ functors from a $𝕜\Bbbk$-linear category with essential image $\mathrm{essim}\left(\psi_{\eta}\right)$ can be computed on $f_{*}\mathrm{Inj}_{f_{*}\eta}$, c.f. Lemma 5.4.
More precisely let
$\varphi=\left(\varphi_{1},...,\varphi_{i}\right):\mathcal{I}\to f_{*}\mathrm{Inj}_{f_{*}\eta}$
be an $\mathcal{A}_{i}$ functor from a $𝕜\Bbbk$-linear category $I$ then we have an isomorphism
$\left(s_{*}\circ h_{*}\circ j\right)_{*}:\mathrm{HH}^{*}\left(\mathcal{I},f_{*}\mathrm{Inj}_{f_{*}\eta}\right)_{2-m}$ $\xrightarrow{\sim}\mathrm{HH}^{*}\left(\mathcal{I},\mathcal{D}^{b}\left(\mathrm{Qcoh}Y\right)\right)$
$o_{i}\left(f_{*}\circ\varphi\right)$ $\mapsto o_{i}\left(s_{*}\circ h_{*}\circ j\circ\varphi\circ\iota\right).$
Furthermore an $\mathcal{A}_{i+1}$-lift $\varphi_{i+1}$ of $s_{*}\circ h_{*}\circ j\circ\varphi$ gives rise to an $\mathcal{A}_{i+1}$-functor $\widehat{\varphi_{i+1}}$ such that
$\varphi_{i+1}=s_{*}\circ h_{*}\circ j\circ\widehat{\varphi}_{i+1}.$
Proof.
As $h$ is a homotopy equivalence and $j$ is the inclusion of a full subcategory we have by naturality of obstructions
$\left(h\circ j\right)_{*}:\mathrm{HH}^{m}\left(\mathrm{Inj},f_{*}\mathrm{Inj}_{f_{*}\eta}\right)$ $\xrightarrow{\sim}\mathrm{HH}^{m}\left(\mathrm{Inj},\mathcal{D}^{b}\left(\mathcal{Y}_{0}\right)\right)$
$o_{m}\left(\varphi\right)$ $\mapsto o_{m}\left(h_{*}\circ j\circ\varphi\iota\right).$
For the compatibility with lifts observe that since $j$ and $h$ are homotopy equivalences on the image of the injectives and so admit quasi-inverses they are immediately compatible with intermediate lifts.
As by Proposition 6.7 the functor $s_{*}$ is the inclusion of the upper triangular matrices and by Proposition 6.8 an equivalence on the negative $\mathrm{Ext}$-groups. We get, since $\mathrm{Inj}$ is a $𝕜\Bbbk$-linear category
$s_{*}:\mathrm{HH}^{m}\left(\mathrm{Inj},\mathcal{D}^{b}\left(\mathcal{Y}_{0}\right)\right)_{<0}$ $\xrightarrow{\sim}\mathrm{HH}^{m}\left(\mathrm{Inj},\mathcal{D}^{b}\left(\mathrm{Qcoh}Y\right)\right)_{<0}$
$o_{m}\left(h_{*}\circ j\circ\varphi\right)$ $\mapsto o_{m}\left(s_{*}\circ h_{*}\circ j\circ\varphi\right).$
Since any intermediate lift would land in the $\mathcal{A}_{\infty}$-submodule of morphisms in negative degrees and $s_{*}$ is an isomorphism on these we have that also this comparison morphism is compatible with intermediate lifts.
So altogether we have the desired isomorphism and compatibility with lifts. ∎
The above Lemma 7.2 has the following immediate consequence
Corollary 7.3.
The functor $\psi_{\eta}\circ\iota:\mathrm{Inj}\to\mathcal{D}^{b}\left(\mathrm{Qcoh}Y\right)$ has an $\mathcal{A}_{\infty}$-lift if and only if $f_{*}\circ L\circ\iota:\mathrm{Inj}\to f_{*}\mathrm{Inj}_{f_{*}\eta}$ has an $\mathcal{A}_{\infty}$-lift.
Now we may restrict for the computation of the obstructions
$o_{i}\left(\psi_{\eta}\circ\iota\right)\in\mathrm{HH}^{i}\left(\mathrm{Inj},\mathrm{H}^{2-i}\mathcal{D}^{b}_{\mathrm{dg}}\left(\mathcal{Y}\right)\right)$
to the image of $f_{*}\circ L$ in negative degrees. In particular, as we consider minimal $\mathcal{A}_{\infty}$-structures, the $\mathrm{Inj}$-bimodule structure on
$\mathrm{H}^{<0}\mathrm{im}\left(f_{*}\circ L\circ\iota\right)$ $\cong\mathrm{Ext}^{*-2+m}\left(f_{*}I,f_{*}\left(I\otimes M\right)\right)$ Corollary 6.2
is given by the $𝕜\Bbbk$-linear functor
$\mathrm{H}^{<0}\mathrm{im}\left(f_{*}\circ L\circ\iota\right):\mathrm{Inj}\otimes\mathrm{Inj}^{\mathrm{op}}$ $𝕜\displaystyle\to\mathcal{D}^{b}_{\infty}\left(\Bbbk\right)$
$\left(I,J\right)$ $\mapsto\mathrm{Ext}^{*-2+m}_{Y}\left(f_{*}I,f_{*}J\otimes M\right)$
$\left(\alpha\otimes\beta\right)$ $\mapsto f_{*}\circ L\left(\alpha\right)\otimes\left(f_{*}\circ L\left(\beta\right)\otimes f_{*}M\right).$
Lemma 7.4.
The functor $f_{*}:\mathrm{Inj}_{\eta}\to f_{*}\mathrm{Inj}_{f_{*}\eta}$ is an equivalence on $\mathrm{H}^{2-m}\mathrm{Inj}_{\eta}$.
$f_{*}:\mathrm{H}^{2-m}\mathrm{Inj}_{\eta}\xrightarrow{\sim}\mathrm{H}^{2-m}f_{*}\mathrm{Inj}_{f_{*}\eta}$
Proof.
We know that $f_{*}$ is by construction literally the functor $f_{*}$ associated to the closed immersion of a divisor $f:X\hookrightarrow Y$ on the two components of
$\mathrm{Inj}_{\eta}\left(I,J\right)\cong\mathrm{Ext}^{0}_{X}\left(I,J\right)\oplus\mathrm{Ext}^{0}_{X}\left(I,J\otimes M\right)[2-m].$
Now, as $f_{*}$ is induced by a direct image along a closed immersion it is an isomorphism on $\mathrm{Ext}^{0}$-groups. In particular
$f_{*}:\mathrm{H}^{0}\mathrm{Inj}_{\eta}$ $\xrightarrow{\sim}\mathrm{H}^{0}f_{*}\mathrm{Inj}_{f_{*}\eta}$
$f_{*}:\mathrm{H}^{2-m}\mathrm{Inj}_{\eta}$ $\xrightarrow{\sim}\mathrm{H}^{2-m}f_{*}\mathrm{Inj}_{f_{*}\eta}$
as claimed. ∎
Proposition 7.5.
The functor $\mathrm{H}^{<0}\mathrm{im}\left(f_{*}\circ L\circ\iota\right)$ is naturally an $\mathrm{Inj}$-$\mathcal{A}_{\infty}$-bimodule with action given via the functor $f_{*}\circ L$.
Proof.
We have that $\mathrm{Inj}$ is $𝕜\Bbbk$-linear and $\mathrm{H}^{<0}\mathrm{im}\left(f_{*}\circ L\circ\iota\right)$ is concentrated in degrees $2-m$ and $3-m$. So the only possible higher multiplications of $\mathrm{Inj}$ on $\mathrm{H}^{<0}\mathrm{im}\left(f_{*}\circ L\circ\iota\right)$ are $\mathrm{m}_{2}$ and $\mathrm{m}_{3}$. By construction we have that $f_{*}\circ L$ is a functor, so it is compatible with $\mathrm{m}_{2}$ and in particular associative, i.e.
$\left[\mathrm{m}^{f_{*}\circ L}_{2},m^{f_{*}\circ L}_{2}\right]=0.$
Now for $\mathrm{m}_{3}$ we need to verify
$\left[\mathrm{m}^{f_{*}\circ L}_{3},\mathrm{m}^{f_{*}\circ L}_{2}\right]=0.$
This is by minimality the $\mathcal{A}_{\infty}$-equation for the whole category
$\mathrm{H}^{*}\mathrm{im}\left(f_{*}\circ L\circ\iota\right)\cong f_{*}\mathrm{Inj}_{f_{*}\eta}.$
In particular
$\left[\mathrm{m}^{f_{*}\mathrm{Inj}_{f_{*}\eta}}_{3},\mathrm{m}^{f_{*}\mathrm{Inj}_{f_{*}\eta}}_{2}\right]=\left[\mathrm{m}^{f_{*}\circ L}_{3},\mathrm{m}^{f_{*}\circ L}_{2}\right]=0$
holds.
Before we continue to the computation of $\mathrm{m}_{3}$ on $\mathrm{H}^{<0}\mathrm{im}\left(f_{*}\circ L\circ\iota\right)$ as an $\mathrm{Inj}$-bimodule we recall the following fact.
Lemma 7.6.
Let $f^{*}\dashv f_{*}:\mathcal{D}_{\infty}\left(\mathrm{Qcoh}X\right)\to\mathcal{D}_{\infty}\left(\mathrm{Qcoh}Y\right)$ be an adjuction of $\mathcal{A}_{\infty}$-functors and let $N$ be an $\mathcal{A}_{\infty}$-$I$-bimodule of the shape
$N:I\otimes I^{\mathrm{op}}$ $\to\mathcal{D}^{b}\left(\mathrm{Qcoh}X\right)$
$\left(i,j\right)$ $\mapsto\mathrm{Ext}^{*}_{X}\left(Ni,Nj\right)$
$\alpha\otimes\beta$ $\mapsto N_{1}\left(\alpha\right)\otimes N_{2}\left(\beta\right)$
for $I$ a $𝕜\Bbbk$-linear category. Then we have that the adjunction quasi-isomorphism
$\psi:\mathrm{Ext}^{*}_{\mathcal{B}}\left(f_{*}X,f_{*}Y\right)\xrightarrow{\sim}\mathrm{Ext}^{*}_{\mathcal{A}}\left(f^{*}f_{*}X,Y\right)$
induces a quasi-isomorphism of $\mathcal{A}_{\infty}$-$I$-bimodules.
$\begin{pmatrix}f_{*}N:I\otimes I^{\mathrm{op}}\to\mathcal{D}^{b}\left(\mathrm{Qcoh}Y\right)\\ \left(i,j\right)\mapsto\mathrm{Ext}^{*}_{X}\left(f_{*}Ni,f_{*}Nj\right)\\ \alpha\otimes\beta\mapsto f_{*}N_{1}\left(\alpha\right)\otimes f_{*}N_{2}\left(\beta\right)\end{pmatrix}\cong\begin{pmatrix}I\otimes I^{\mathrm{op}}\to\mathcal{D}^{b}\left(\mathrm{Qcoh}X\right)\\ \left(i,j\right)\mapsto\mathrm{Ext}^{*}_{X}\left(f^{*}f_{*}Ni,Nj\right)\\ \alpha\otimes\beta\mapsto f^{*}f_{*}N_{1}\left(\alpha\right)\otimes N_{2}\left(\beta\right)\end{pmatrix}.$
Proof.
Consider a $\mathrm{dg}$-replacement of the adjunction
$f^{*}\dashv f_{*}:\mathcal{D}_{\mathrm{dg}}\left(\mathrm{Qcoh}X\right)\to\mathcal{D}_{\mathrm{dg}}\left(\mathrm{Qcoh}Y\right).$
Then it suffices to show that the isomorphism
$\mathcal{D}_{\mathrm{dg}}\left(\mathrm{Qcoh}Y\right)\left(f_{*}Ni,f_{*}Nj\right)\xrightarrow{\sim}\mathcal{D}_{\mathrm{dg}}\left(\mathrm{Qcoh}X\right)\left(f^{*}f_{*}Ni,Nj\right)$
is a quasi-isomorphism of $\mathrm{dg}$-$I$-bimodules. By Definition of $\mathrm{dg}$-adjunction it is already a quasi-isomorphism of chain complexes of vectorspaces. In particular it suffices to prove that it is a morphism of $I$-bimodules.
For this consider first the operation from the left on $f_{*}N$. Denote it by
$f_{*}N_{1}\left(\alpha\right):f_{*}Ni^{\prime}\to f_{*}Ni$
, then we have the commutative diagram
[…]
Now for the action on the left given by $f_{*}N_{2}\left(\beta\right):f_{*}Nj\to f_{*}Nj^{\prime}$ we get analogously
[…]
In particular
$\mathcal{D}_{\mathrm{dg}}\left(\mathrm{Qcoh}\left(Y\right)\right)\left(f_{*}Ni,f_{*}Nj\right)\xrightarrow{\sim}\mathcal{D}_{\mathrm{dg}}\left(\mathrm{Qcoh}X\right)\left(f^{*}f_{*}Ni,Nj\right)$
is compatible with both actions and induces a quasi-isomorphism of $\mathcal{A}_{\infty}$-$I$-bimodules. ∎
Remark 7.7.
In the above proof it is essential that we only consider the bimodule structure given by the functor $f_{*}$. Otherwise we would not be able to compare the right action, as not every functor in the essential image arises via $f_{*}$.
The following is a rewritten proof to our setup of the well known fact that for a divisor $X\hookrightarrow Y$ the object $f_{*}\mathcal{F}\in\mathcal{D}_{\infty}\left(Y\right)$ usually is not formal over $Y$ for a sheaf $\mathcal{F}$, however, it is formal over $X$. In particular it is essential that we restrict in this proof to the bimodule structure over $\mathrm{Inj}$.
Proposition 7.8.
The $\mathrm{Inj}$-$\mathcal{A}_{\infty}$-bimodule $\mathrm{H}^{<0}\mathrm{im}\left(f_{*}\circ L\circ\iota\right)$ has $\mathrm{m}_{3}=0$.
Proof.
By adjunction we have a natural $\mathcal{A}_{\infty}$-bimodule-isomorphism written down objectwise
$\mathrm{Ext}_{Y}^{*-2+m}\left(f_{*}I,f_{*}J\otimes M\right)\cong\mathrm{Ext}^{*-2+m}_{X}\left(f^{*}f_{*}I,J\otimes M\right).$
This isomorphism is by Lemma 7.6 compatible with the bimodule structure induced by $\mathcal{X}$ on the $\mathcal{X}$-bimodule $\mathrm{H}^{<0}\mathrm{im}\left(f_{*}\circ L\circ\iota\right)$ and so in particular compatible with the $\mathrm{Inj}$-bimodule structure induced by $f_{*}\circ L\circ\iota$ from Proposition 7.5.
So we have the following chain of isomorphisms of $\mathrm{Inj}$-$\mathcal{A}_{\infty}$-bimodules, written down objectwise:
$\mathrm{H}^{<0}\mathrm{im}\left(f_{*}\circ L\circ\iota\right)\left(I,J\right)$
$\cong\mathrm{Ext}^{*-2+m}_{X}\left(f^{*}f_{*}I,J\otimes M\right)$ $f^{*}\vdash f_{*}$
$\cong\mathrm{Ext}^{*-2+m}_{X}\left(f^{*}f_{*}\mathcal{O}_{X}\otimes_{\mathcal{O}_{Y}}I,J\otimes M\right)$
$\cong\mathrm{Ext}^{*-2+m}_{X}\left(I\otimes C_{f}[1]\oplus I,J\otimes M\right)$ $f^{*}f_{*}\mathcal{O}_{X}\cong\otimes\left(\mathcal{O}_{X}\oplus C_{f}[1]\right)$
$\cong\mathrm{Ext}^{*-2+m}_{X}\left(I,J\otimes M\right)\oplus\mathrm{Ext}^{*-3+m}_{X}\left(I\otimes C_{f},J\otimes M\right)$ $\mathrm{Ext}^{*}\text{ is additive}$
where all isomorphisms hold on the $\mathcal{A}_{\infty}$-level. The first one is an isomorphisms of $\mathcal{A}_{\infty}$-$\mathrm{Inj}$-bimodules by Lemma 7.6, the second is the defintition of the pullback, the third one is $\mathcal{A}_{\infty}$/$\mathrm{dg}$ by [AC12, 0.7] and the last one is $\mathcal{A}_{\infty}$ as it is induced by the additivity of $\mathrm{Ext}$. Altogether this sequence of natural isomorphisms operates on actions via:
$f_{*}\circ L\left(\alpha\right)$ $\mapsto\begin{pmatrix}L\left(\alpha\right)\\ L\left(\alpha\right)\otimes C_{f}\end{pmatrix}$
$f_{*}\circ L\left(\beta\right)$ $\mapsto\begin{pmatrix}L\left(\alpha\right)\\ L\left(\alpha\right)\end{pmatrix}.$
As $\mathrm{Inj}$ is $𝕜\Bbbk$-linear and both direct summands are concentrated in one degree we have that the $\mathrm{m}_{3}$ on $\mathrm{Ext}^{*}_{X}\left(I,J\otimes M\right)$ and $\mathrm{Ext}^{*}_{X}\left(I\left(-X\right),J\otimes M\right)$ vanishes. In particular we found a minimal $\mathcal{A}_{\infty}$-structure with vanishing $\mathrm{m}_{3}$ of the $\mathrm{Inj}$-$\mathcal{A}_{\infty}$-bimodule
$\mathrm{H}^{<0}\mathrm{im}\left(f_{*}\circ L\circ\iota\right)\cong\mathrm{H}^{3-m}\mathrm{im}\left(f_{*}\circ L\circ\iota\right)\oplus\mathrm{H}^{2-m}\mathrm{im}\left(f_{*}\circ L\circ\iota\right)$
as claimed. ∎
Remark 7.9.
Even though we just proved that $\mathrm{m}_{3}=0$ on the $\mathcal{A}_{\infty}$-bimodule $\mathrm{H}^{<0}\mathrm{im}f_{*}\circ L\circ\iota$, this does not apply to $\mathrm{m}_{m}=f_{*}\eta$ on $\mathrm{H}^{*}\mathrm{im}f_{*}\circ L\circ\iota$. Even more, due to $\mathrm{m}_{\geq 3}$ vanishing on the $\mathrm{Inj}$-bimodule $\mathrm{H}^{\leq 0}\mathrm{im}f_{*}\circ L\circ\iota$, we actually have that $\mathrm{im}f_{*}\circ L$ has a non-vanishing $\mathrm{m}_{m}=f_{*}\eta$ as an $\mathrm{Inj}$-bimodule.
Lemma 7.10.
We have independent of choices
$0\neq o_{m}\left(f_{*}\circ L\circ\iota\right)\in\mathrm{HH}^{*}\left(\mathrm{Inj},\mathrm{H}^{<0}\mathrm{\mathrm{im}}f_{*}\circ L\circ\iota\right)$
Proof.
By naturality of obstructions we have
$o_{m-1}\left(f_{*}\circ L\circ\iota\right)$ $=f_{*}o_{m-1}\left(L\circ\iota\right)=0$
In particular we may choose an $\mathcal{A}_{m-1}$ lift
$\left(f_{*}\circ L\circ\iota,0,...,0,\delta,g\right):\mathrm{Inj}\to\mathrm{H}^{\leq 0}\mathrm{\mathrm{im}}f_{*}\circ L\circ\iota\cong f_{*}\mathrm{Inj}_{f_{*}\eta}.$
By Proposition 7.8 we have
$\mathrm{m}_{3}=0$
on the $\mathrm{Inj}$-bimodule $\mathrm{H}^{<0}\mathrm{\mathrm{im}}f_{*}\circ L\circ\iota$. Now we compute in $\mathrm{HH}^{m}\left(\mathrm{Inj},\mathrm{H}^{2-m}\mathrm{\mathrm{im}}\psi_{\eta}\circ\iota\right)$ that the obstruction is independent of choices
[…]
$=[\mathrm{m}_{m},f_{*}\circ L\circ\iota]\pm[\mathrm{m}_{3},\delta]$ $[\mathrm{m}_{2},\_]=\mathrm{d}_{\mathrm{Hoch}}$
$=[\mathrm{m}_{m},f_{*}\circ L\circ\iota]$ Proposition 7.8
$=\mathrm{m}_{m}\left(f_{*}\circ L\circ\iota,...,f_{*}\circ L\circ\iota\right)$ Inj has $\mathrm{Inj}\text{ has }\mathrm{m}_{m}=0$
$=f_{*}\mathrm{m}_{m}\left(L\circ\iota,...,L\circ\iota\right)$ $f_{*}\text{ is }\mathcal{A}_{\infty}$
$=f_{*}o_{m}\left(L\circ\iota\right)$
$\neq 0.$
Where we used in the last line that $f_{*}$ induces an equivalence by Lemma 7.4 on
$\mathrm{Ext}_{\mathcal{X}_{\eta}}^{2-m}\left(L\left(I\right),L\left(J\right)\right)\cong\mathrm{Ext}_{\mathcal{Y}_{f_{*}\eta}}^{2-m}\left(f_{*}\circ L\left(I\right),f_{*}\circ L\left(J\right)\right)$
In particular the obstruction does not vanish and is independent of choices. ∎
Proposition 7.11.
We have independent of choices
$0\neq o_{m}\left(\psi_{\eta}\circ\iota\right).$
Proof.
By Lemma 7.2 we have have
$o_{m}\left(\psi_{\eta}\circ\iota\right)=o_{m}\left(f_{*}\circ L\circ\iota\right)$
and by Lemma 7.10
$o_{m}\left(f_{*}\circ L\circ\iota\right)\neq 0.$
Which finishes the claim. ∎
Theorem 7.12.
The functor $\psi_{\eta}$ does not admit an $\mathcal{A}_{\infty}$/$\mathrm{dg}$-lift.
Proof.
By Lemma 7.11 we have that $o_{m}\left(\psi_{\eta}\circ\iota\right)$ does not vanish and is independent of choices. In particular $\psi_{\eta}\circ\iota$ cannot have an $\mathcal{A}_{\infty}$-lift. Since $\iota$ is dg, this means that $\psi_{\eta}$ cannot have an $\mathcal{A}_{\infty}$-lift. ∎
By [Kü22, Theorem 4.14] we can compute for a smooth hypersurface the kernels of $\mathrm{HH}^{*}\left(X,\mathcal{O}\left(p\right)\right)$ explicitly via twisted Hodge-diamonds, in particular all these kernels in $\mathrm{HH}^{\geq\dim X+3}\left(X,\mathcal{O}\left(p\right)\right)$ define non-$\mathrm{dg}$ functors.
Corollary 7.13.
Let $f:X\hookrightarrow\mathbb{P}^{n+1}$ be a smooth degree $d$-hypersurface. Then we have that the space of non-Fourier-Mukai functors arising as deformation of the shape $\psi_{\eta}$ can be computed explicitly .
We now give a few examples of Twisted Hodge diamonds and the resulting dimensions for non-$\mathrm{dg}$ deformations of the closed immersions we computed using the Sage package [BG].
Example 7.14.
Let $f:X\hookrightarrow\mathbb{P}^{7}$ be a smooth degree $7$ hypersurface then its Hodge diamond is
$\begin{matrix}\;&\;&\;&\;&\;&\;&\;&1&\;&\;&\;&\;&\;\\ \;&\;&\;&\;&\;&\;&0&\;&0&\;&\;&\;&\;\\ \;&\;&\;&\;&\;&0&\;&1&\;&0&\;&\;&\;\\ \;&\;&\;&\;&0&\;&0&\;&0&\;&0&\;&\;\\ \;&\;&\;&0&\;&0&\;&1&\;&0&\;&0&\;\\ \;&\;&0&\;&0&\;&0&\;&0&\;&0&\;&0\\ \;&0&\;&\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\mathclap{36}\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}&\;&\mathclap{2472}&\;&\mathclap{8093}&\;&\mathclap{2472}&\;&\mathclap{36}&\;&0\\ \;&\;&0&\;&0&\;&0&\;&0&\;&0&\;&0\\ \;&\;&\;&0&\;&0&\;&1&\;&0&\;&0&\;&\;&\;\\ \;&\;&\;&\;&0&\;&0&\;&0&\;&0&\;&\;&\;&\;\\ \;&\;&\;&\;&\;&0&\;&1&\;&0&\;&\;&\;&\;&\;\\ \;&\;&\;&\;&\;&\;&0&\;&0&\;&\;&\;&\;&\;&\;\\ \;&\;&\;&\;&\;&\;&\;&1.&\;&\;&\;&\;&\;&\;&\;\\ \end{matrix}$
And so we have by [Kü22, Theorem 4.14],
$\dim\ker\left(f_{*}:\mathrm{HH}^{10}\left(X,\mathcal{O}_{X}\left(1\right)\right)\to\mathrm{HH}^{10}\left(\mathbb{P}^{7},f_{*}\mathcal{O}_{X}\left(1\right)\right)\right)$ $=36$
All non-zero $\eta$ in this $36$-dimensional space give rise to non-Fourier-Mukai functors.
Example 7.15.
Let $f:X\hookrightarrow\mathbb{P}^{7}$ be a smooth degree $7$ hypersurface then its $8$ twisted Hodge diamond is
$\begin{matrix}\;&\;&\;&\;&\;&0&\;&\;&\;&\;&\;\\ \;&\;&\;&\;&0&\;&0&\;&\;&\;&\;\\ \;&\;&\;&0&\;&0&\;&0&\;&\;&\;\\ \;&\;&0&\;&0&\;&0&\;&0&\;&\;\\ \;&0&\;&0&\;&0&\;&0&\;&0&\;\\ \mathllap{299}6&\;&\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\mathclap{20993}\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}&\;&\mathclap{15267}&\;&\mathclap{917}&\;&0&\;&0\\ \;&\mathllap{157}5&\;&0&\;&0&\;&0&\;&0&\;\\ \;&\;&\mathllap{577}5&\;&0&\;&0&\;&0&\;&\;\\ \;&\;&\;&\mathllap{1039}5&\;&0&\;&0&\;&\;&\;\\ \;&\;&\;&\;&\mathllap{900}2&\;&0&\;&\;&\;&\;\\ \;&\;&\;&\;&\;&\mathllap{299}6.&\;&\;&\;&\;&\;\\ \end{matrix}$
And so we have by [Kü22, Theorem 4.14],
$\dim\ker\left(f_{*}:\mathrm{HH}^{8}\left(X,\mathcal{O}_{X}\left(-8\right)\right)\to\mathrm{HH}^{8}\left(X,f_{*}\mathcal{O}_{X}\left(-8\right)\right)\right)$ $=\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}20993\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}.$
Again every single $0\neq\eta$ in this $20993$-dimensional space give rise to a non-Fourier-Mukai functors.
Finally, paired with [Kü22, Proposition 4.16] we get the following corollary
Corollary 7.16.
Let $f:X\hookrightarrow\mathbb{P}^{n+1}$ be a smooth degree $d$-hypersurface. Then there exists a non-Fourier-Mukai functor
$\psi_{\eta}:\mathcal{D}^{b}\left(\mathrm{Qcoh}X\right)\to\mathcal{D}^{b}\left(\mathrm{Qcoh}\mathbb{P}^{n+1}\right).$
Finally recall that all the above constructions are not liftable in the spectral setting as well by the following proposition which is the main reason we considered a field of characteristic zero.
Proposition 7.17.
[RVdBN19, Proposition B.2.1.] If $𝕜\Bbbk=\mathbb{Q}$ then the functor $\psi_{\eta}$ does not admit a spectral lift.

### v2

@@
 Abstract.
+We use injectives as a big tilting object to obstruct liftability of exact functors to the $\mathrm{dg}$-level.
 We use the inclusion of injectives into the canonical heart as a replacement for tilting objects in computations of the characteristic morphism.
@@
 1. Introduction
-The idea behind Fourier-Mukai functors as categorified correspondences was introduced by S.
+The idea behind Fourier–Mukai functors as categorified correspondences was introduced by S.
 Mukai [Muk81] to understand the relation between the derived category of an abelian variety $\mathcal{A}$ and its dual $\mathcal{A}^{\vee}$.
-Its classical definition being:
+The classical definition is:
 Definition 1.1.
@@
 $f\colon\mathcal{D}^{b}\left(\mathrm{Qcoh}X\right)\to\mathcal{D}^{b}\left(\mathrm{Qcoh}Y\right),$
-for two schemes $X$ and $Y$ is Fourier-Mukai if there exists an $M$ in $\mathcal{D}^{b}\left(\mathrm{coh}\left(X\times Y\right)\right)$, respectively in $\mathcal{D}^{b}\left(\mathrm{Qcoh}\left(X\times Y\right)\right)$, such that
+for two schemes $X$ and $Y$ is Fourier-Mukai if there exists a sheaf $M$ in $\mathcal{D}^{b}\left(\mathrm{coh}\left(X\times Y\right)\right)$, respectively in $\mathcal{D}^{b}\left(\mathrm{Qcoh}\left(X\times Y\right)\right)$, such that
 $f\cong\pi_{1,*}\left(M\otimes\pi_{2}^{*}\_\right).$
-This concept allows the study of functors from a geometric point of view and in particular the application of geometric arguments and constructions to them.
+This concept allows for the study of functors from a geometric point of view and in particular the application of geometric arguments and constructions to them.
 With time it was realized that in fact every known functor between derived categories of smooth projective schemes is Fourier-Mukai.
-For example push-forward and pull-back along a morphism of schemes $f$ is induced by taking $M$ to be the structure sheaf of its graph $\mathcal{O}_{\Gamma_{f}}$.
-A celebrated result by D. Orlov gave a criterion to check whether a functor is Fourier-Mukai,
+For example, push-forward and pull-back along a morphism of schemes $f$ is induced by taking $M$ to be the structure sheaf of its graph $\mathcal{O}_{\Gamma_{f}}$.
+A celebrated result by D. Orlov gave a criterion to check whether a functor is Fourier-Mukai:
 Theorem 1.2.
-[Orl97] Let $f:\mathcal{D}^{b}\left(\mathrm{coh}X\right)\to\mathcal{D}^{b}\left(\mathrm{coh}Y\right)$ be a fully-faithful functor between derived categories of smooth projective schemes.
+[Orl97] Let $f:\mathcal{D}^{b}\left(\mathrm{coh}X\right)\to\mathcal{D}^{b}\left(\mathrm{coh}Y\right)$ be a fullyfaithful functor between derived categories of smooth projective schemes.
 Then $f$ is Fourier-Mukai.
-Whereas B. Toën realized a beautiful conceptual categorization of the property,
+Whereas B. Toën realized a beautiful conceptual categorization of the property.
 Theorem 1.3.
@@
 $\widehat{f}:\mathcal{D}_{\mathrm{dg}}^{b}\left(\mathrm{Qcoh}X\right)\to\mathcal{D}_{\mathrm{dg}}^{b}\left(\mathrm{Qcoh}Y\right),$
-where $\mathcal{D}_{\mathrm{dg}}^{b}\left(\_\right)$ is a $\mathrm{dg}$-enhancement of $\mathcal{D}^{b}\left(\_\right)$.
+such that $\mathrm{H}^{0}\left(\widehat{f}\right)\cong f$, where $\mathcal{D}_{\mathrm{dg}}^{b}\left(\_\right)$ is a $\mathrm{dg}$-enhancement of $\mathcal{D}^{b}\left(\_\right)$.
 This theorem in particular allows the interpretation of Fourier-Mukai functors as the geometric analog to bimodules and their represented functors.
-These observations led to a conjecture that actually all exact functors between derived categories of smooth projective schemes are Fourier-Mukai.
+These observations led to a conjecture that all exact functors between derived categories of smooth projective schemes are Fourier-Mukai.
 Conjecture 1.4.
@@
 that is not Fourier-Mukai, where $Q_{3}\hookrightarrow\mathbb{P}^{4}$ is a smooth three dimensional quadric.
-We were able to generalize their argument to higher dimensions in [Kü22] using twisted Hodge diamonds, while T.
+We were able to generalize their argument to higher dimensions in [Kü24] using twisted Hodge diamonds, while T.
 Raedschaelders, A.
@@
 Van den Bergh [RRVdB19] were able to construct similar non-Fourier-Mukai functors by losing control over the target space.
-A big drawback of all these approaches is that they heavily rely on the existence of tilting objects in the source category in order to control liftability.
+A big drawback of all these approaches is that they heavily rely on the existence of tilting objects in the source category in order to control liftability, which tend to be rare.
 We avoid this issue here by replacing the tilting bundle with the subcategory of injectives in the canonical heart of the derived category of quasi-coherent sheaves.
@@
 Corollary 8.13.
-Let $f:X\hookrightarrow\mathbb{P}^{n+1}$ be a smooth degree $d$-hypersurface. Then the space of non-Fourier-Mukai functors arising as deformation of the shape $\psi_{\eta}$ can be computed explicitly.
+Let $f:X\hookrightarrow\mathbb{P}^{n+1}$ be a smooth degree $d$-hypersurface.
+Then the space of non-Fourier-Mukai functors arising as deformations of the shape $\psi_{\eta}$ can be computed explicitly.
+Finally we show a possible avenue for generalizing our result to coherent sheaves by using extension results by J. Lurie [Lur09] respectively B. Toën [Toe07].
+Theorem 8.12.
+Assume that $\psi_{\eta}$ is continuous. Then the restriction
+$\psi_{\eta}:\mathcal{D}^{b}\left(\mathrm{coh}X\right)\to\mathcal{D}^{b}\left(\mathrm{coh}Y\right)$
+is not Fourier-Mukai.
 1.1. Proof strategy
@@
 The key insight in our approach is that the non-liftability of $\psi_{\eta}\left(T\right)$ boils down to the the functor $\psi_{\eta}\circ\iota_{T}:*_{\Gamma}\to\mathcal{D}^{b}\left(Y\right)$ not admitting an $\mathcal{A}_{\infty}$-lift, here we denoted by $*_{\Gamma}$ the $𝕜\Bbbk$-linear category with one object * and endomorphisms $\Gamma$.
-In particular, since $T$ was assumed to be tilting the functor
+In particular, since $T$ was assumed to be tilting, the functor
 $\iota_{T}:*_{\Gamma}\to\mathcal{D}^{b}\left(X\right)$
@@
 By classical homological algebra this functor completely controls $\mathcal{D}^{b}\left(\mathrm{Qcoh}X\right)$ [Wei94].
-In order to explain our proof strategry, consider the follwoing diagram
+In order to explain our proof strategy, consider the following diagram
 $\mathrm{Inj}$$\mathcal{D}^{b}\left(\mathrm{Qcoh}X\right)$$\mathcal{D}^{b}\left(\mathcal{X}_{\eta}\right)$$\mathcal{D}^{b}\left(\mathrm{Qcoh}Y\right),$$\mathcal{D}_{\mathrm{dg}}^{b}\left(\mathrm{Qcoh}X\right)$$\mathcal{D}_{\mathrm{dg}}^{b}\left(\mathcal{X}_{\eta}\right)$$\mathcal{D}_{\mathrm{dg}}^{b}\left(\mathrm{Qcoh}Y\right)$.$\iota$$\iota$$L$$\widetilde{f}_{*}$$\widetilde{f}_{*}$
@@
 The plan is to push this obstruction forward.
-In contrast to [RVdBN19] and [Kü22] this needs to be done with a lot of care as we do not have access to Serre duality in our setting.
+In contrast to [RVdBN19] and [Kü24] this needs to be done with a lot of care as we do not have access to Serre duality in our setting.
 In particular we study the concrete construction of $\widetilde{f}_{*}$ to compute that the obstruction $o_{m}\left(\psi_{\eta}\circ\iota\right)$ indeed is independent of choices and coincides with the pushed forward obstruction $f_{*}o_{m}\left(L\circ\iota\right)$.
@@
 1.2. Structure of the paper
-In Section 2.1 we recall preliminaries on Hochschild coholomology and the characteristic morphism in the classical setting.
+In Section 2.1 we recall preliminaries on Hochschild cohomology and the characteristic morphism in the classical setting.
 In particular we recall in Definition 2.5 the construction of an infinitesimal deformation along a Hochschild cohomology class, which is a key player in the construction of $\psi_{\eta}.$
@@
 In Section 5 we then collect the properties, definitions and computations of $\mathcal{A}_{\infty}$-obstructions that we will use later.
-Since we are not using tilting objects in the category of coherent sheaves we are not able to use Serre duality as in [RVdBN19, Kü22] in order to show that the relevant $\mathrm{Ext}$-groups are concentrated in degree $-1$.
+Since we are not using tilting objects in the category of coherent sheaves we are not able to use Serre duality as in [RVdBN19, Kü24] in order to show that the relevant $\mathrm{Ext}$-groups are concentrated in degree $-1$.
 So we will have to control the $\mathcal{A}_{\infty}$-structure on our target category more explicitly.
 For this we study in Section 6 the explicit structure of the functor $\widetilde{f}$.
-In particular we completely understand the essential image and its $\mathcal{A}_{\infty}$-bimodule structure.
-Then we finish in Section 7 by computing explicitly that the obstructions of $\psi_{\eta}$ cannot vanish.
-We then combine this result with computations from [Kü22] in order to conclude the existence of a vast amount of non-Fourier-Mukai functors with well-behaved target categories.
+In particular we completely understand its essential image and its $\mathcal{A}_{\infty}$-bimodule structure.
+We finish in Section 7 by computing explicitly that the obstructions of $\psi_{\eta}$ cannot vanish.
+We then combine this result with computations from [Kü24] in order to conclude the existence of a vast amount of non-Fourier-Mukai functors with well-behaved target categories and showing possible arguments for the case of coherent sheaves.
 Acknowledgements
@@
 Špenko for her awesome support and helpful discussions.
-Finally I would like to thank B.
+I would like to thank B.
 Briggs and A.
 Cǎldǎraru for their insights on Hochschild cohomology, $\mathcal{A}_{\infty}$-structures and formality of sheaves over source spaces.
+Finally I are deeply grateful for the insight of J.
+Holstein on continuation of functors allowing the possible restriction to coherent sheaves.
 Notation
@@
 We consider all schemes and categories to be over $𝕜\Bbbk$.
-For a $𝕜\Bbbk$-linear category $\mathcal{X}$, denote by $\mathcal{D}^{b}\left(\mathcal{X}\right)$ the bounded derived category of modules over $\mathcal{X}$.
-We will consider left modules.
+For a $𝕜\Bbbk$-linear category $\mathcal{X}$, denote by $\mathcal{D}^{b}\left(\mathcal{X}\right)$ the bounded derived category of left modules over $\mathcal{X}$.
+For $X$ a scheme we denote by $\mathcal{D}^{b}\left(X\right)$ the bounded derived category of coherent sheaves.
 Denote by $\mathcal{D}_{\mathrm{dg}}^{b}\left(\_\right)$ the (canonical) dg-enhancement of $\mathcal{D}^{b}\left(\_\right)$.
 We denote by $\mathcal{D}^{b}_{\mathcal{C}}\left(\_\right)$ the full subcategory of $\mathcal{D}^{b}\left(\_\right)$ with cohomology objects in $\mathcal{C}$.
-We will also refer to $\mathrm{dg}$ respectively $\mathcal{A}_{\infty}$-functors between smooth geometric categories as Fourier Mukai functors following the insight from B.
+We also use the slightly strange convention of $f_{*}$ for the pullback of $𝕜\Bbbk$-linear categories in order to mimic the geometric setting and improve readability.
+We will also refer to $\mathrm{dg}$ respectively $\mathcal{A}_{\infty}$-functors between smooth geometric categories as Fourier Mukai functors following the insight of B.
 Toën [Toe07].
-We also use the shorthand
+Finally we use the shorthand
 $\left[f_{i},m_{k}\right]:=$ $\sum_{\mathclap{r+t=i-1}}\left(-1\right)^{r+kt}f_{i}\left(\mathrm{Id}^{\otimes r}\otimes\mathrm{m}_{k}\otimes\mathrm{Id}^{\otimes t}\right)$
 $-\sum_{\mathclap{1\leq j\leq k}}\left(-1\right)^{ij}\mathrm{m}_{k}\left(\alpha_{1}\otimes...\otimes f_{i}\left(\alpha_{j+1},...,\alpha_{j+i}\right)\otimes...\otimes\alpha_{k+i}\right)$
-for a morphism $\mathcal{A}^{\otimes i}\to\mathcal{M}$ and $\mathcal{M}$ an $\mathcal{A}_{\infty}$-$\mathcal{A}$-module.
+for the Gerstenhaber bracket of a morphism $\mathcal{A}^{\otimes i}\to\mathcal{M}$ and $\mathcal{M}$ an $\mathcal{A}_{\infty}$-$\mathcal{A}$-module.
 2. Preliminaries
-We start off by recalling some basic notions needed for our constructions and arguments.
-For the basic notions on $\mathcal{A}_{\infty}$-structures we refer to the notes by B.
+We start by recalling some basic notions needed for our constructions and arguments.
+For the basics on $\mathcal{A}_{\infty}$-structures we refer to the notes by B.
 Keller [Kel01], the work of H.
@@
 For the construction of $\mathrm{dg}$-hulls we refer to [KKS09].
-2.1. Preliminaries Hochschild cohomology and characteristic morphism
-As we will work a lot with obstructions and Hochschild cohomology we recall the basic constructions and notions here.
+2.1. Hochschild cohomology and characteristic morphism
+As we will work a lot with obstructions in Hochschild cohomology we recall the basic constructions and notions here.
 For a more in depth discussion on the incarnations of Hochschild cohomology we use here, we refer the reader to the original work of G.
@@
 with differential given by
-$df\left(x_{1}\otimes\cdot\otimes x_{n+1}\right):=$ $x_{1}f\left(x_{2}\otimes\cdots\otimes x_{n+1}\right)$
-$+\sum_{i=1}^{n}\left(-1\right)^{i}f\left(x_{1}\otimes\cdots\otimes x_{1}x_{i+1}\otimes\cdots\otimes x_{n+1}\right)$
+$df\left(x_{1}\otimes\cdots\otimes x_{n+1}\right):=$ $x_{1}f\left(x_{2}\otimes\cdots\otimes x_{n+1}\right)$
+$+\sum_{i=1}^{n}\left(-1\right)^{i}f\left(x_{1}\otimes\cdots\otimes x_{i}x_{i+1}\otimes\cdots\otimes x_{n+1}\right)$
 $+\left(-1\right)^{n+1}f\left(x_{1}\otimes\cdots\otimes x_{n}\right)x_{n+1}.$
-The Hochschild cohomology $\mathrm{HH}^{*}\left(\mathcal{X},\mathcal{M}\right)$ is the cohomology of $\mathrm{CC}^{*}\left(\mathcal{X},\mathcal{M}\right)$.
+The Hochschild cohomology $\mathrm{HH}^{*}\left(\mathcal{X},\mathcal{M}\right)$ is the cohomology of $\mathrm{CC}^{\cdot}\left(\mathcal{X},\mathcal{M}\right)$.
 The following classical Lemma 2.2 allows for a more streamlined construction and will help us interpret the Hochschild cohomology in more generality later.
-In particular it highlights the importance of $\mathcal{M}$ being a bimodule.
 Lemma 2.2.
@@
 In bigger generality constructing tensor products of $\mathcal{A}_{\infty}$-categories needs to be done with a lot of care [LRG20].
-However, we will only consider tensor products with $𝕜\Bbbk$-linear categories, which are a lot better behaved and allow our naive approach to work out.
+However, we will only consider tensor products with $𝕜\Bbbk$-linear categories, which are better behaved and allow our naive approach to work.
 The restriction in the following definition to $m\geq 3$ is not necessary in general.
-However, as we will have naturally this assumption and it makes the formulas a lot more concise we present it in this special case.
+However, as we will have naturally this assumption and it makes the formulas more concise we present this special case.
 Definition 2.5.
@@
 $\mathcal{X}_{\eta}\left(x,y\right):=\mathcal{X}\left(x,y\right)\oplus\mathcal{M}\left(x,y\right)\left[2-m\right]$
-and actions given by
+and non-zero multiplications given by
 $\mathrm{m}_{2}:\left(\mathcal{X}\left(x,y\right)\oplus\mathcal{M}\left(x,y\right)\left[2-m\right]\right)^{\otimes 2}$ $\to\mathcal{X}\left(x,y\right)\oplus\mathcal{M}\left(x,y\right)\left[2-m\right]$
-$\left(x,m\right)\otimes\left(x^{\prime},m^{\prime}\right)$ $\mapsto\left(xx^{\prime},xm+m^{\prime}x\right)$
+$\left(x,m\right)\otimes\left(x^{\prime},m^{\prime}\right)$ $\mapsto\left(xx^{\prime},xm^{\prime}+mx^{\prime}\right)$
 $\mathrm{m}_{m}:\left(\mathcal{X}\left(x,y\right)\oplus\mathcal{M}\left(x,y\right)\left[2-m\right]\right)^{\otimes m}$ $\to\mathcal{X}\left(x,y\right)\oplus\mathcal{M}\left(x,y\right)\left[2-m\right]$
-$x_{1}\otimes...\otimes x_{m}$ $\mapsto\eta\left(x_{1},...,x_{m}\right)$
+$x_{1}\otimes...\otimes x_{m}$ $\mapsto\eta\left(x_{1},...,x_{m}\right).$
 This construction comes with a canonical morphism that is the identity on $\mathcal{X}$ and zero on $\mathcal{M}$,
-(1) $\mathcal{X}_{\eta}\to\mathcal{X}.$
+(1) $\mathrm{can}:\mathcal{X}_{\eta}\to\mathcal{X}.$
 Proposition 2.6.
@@
 2.2. Equivariant objects over categories
-In this section we recall the basic notions of equivariant objects and how to interpret them as functors from a small category, which we will apply in the next section to compute the equivariant characteristic morphisms of these objects.
+In this section we recall the basic notions of equivariant objects and how to interpret them as functors from a small category.
+We will apply this in the next section to compute the equivariant characteristic morphisms of these objects.
 We start by recalling the basic definition.
@@
 $\varphi_{Y}\left(\gamma\right)\circ f=f\circ\varphi_{X}\left(\gamma\right).$
-There also exists a $\Gamma$-equivariant incarnation of the characteristic morphism, given by:
+If there is a $\Gamma$-action on $T$ the characteristic morphism from Definition 2.8 admits an equivariant incarnation keeping track of the $\Gamma$-action.
 Definition 2.10.
 Let $\mathcal{X}$ be a $𝕜\Bbbk$-linear category, $\mathcal{M}$ a $𝕜\Bbbk$-central bimodule and let $\mathcal{G}$ be a $\Gamma$-equivariant $\mathcal{X}$-module.
-Then _$\_\otimes\mathcal{G}$ induces the the $\Gamma$-equivariant characteristic morphism
+Then _$\_\otimes\mathcal{G}$ induces the $\Gamma$-equivariant characteristic morphism
 $c_{\Gamma,\mathcal{G}}:\mathrm{HH}^{*}\left(\mathcal{X},\mathcal{M}\right)\to\mathrm{Ext}^{*}_{\mathcal{X}\otimes\Gamma^{\mathrm{op}}}\left(\mathcal{G},\mathcal{M}\otimes\mathcal{G}\right).$
-While $\mathrm{RHom}\left(\_,\mathcal{G}\right)$ induces the the $\Gamma$-equivariant dual characteristic morphism
+While $\mathrm{RHom}\left(\_,\mathcal{G}\right)$ induces the $\Gamma$-equivariant dual characteristic morphism
 $c^{*}_{\Gamma,\mathcal{G}}:\mathrm{HH}^{*}\left(\mathcal{X},\mathcal{M}\right)\to\mathrm{Ext}^{*}_{\mathcal{X}\otimes\Gamma^{\mathrm{op}}}\left(\mathrm{RHom}\left(\mathcal{M},\mathcal{G}\right),\mathcal{G}\right).$
@@
 $\iota_{X,\Gamma}:\mathrm{End}\left(*\right)=\Gamma\longrightarrow X=\iota_{X,\Gamma}.$
-We can do the same construction the other way around, i.e.
-make a $\Gamma$-equivariant object $X$ into a functor via fixing the image of * to be $X$ and the structure morphism the map on morphism spaces.
+Furthermore we can make a $\Gamma$-equivariant object $X$ into a functor via fixing the image of * to be $X$ and the structure morphism given by the map on morphism spaces.
+This gives us two inverse constructions.
 Finally, the naturality requirement for a morphism $\alpha:\iota_{X,\Gamma}\to\iota_{Y,\Gamma}$ in the functor category boils down to
@@
 $\mathcal{I}\to\mathcal{C}.$
-We use the following notation for the category of equivariant objects over $I$ in $\mathcal{C}$ by
+We in particular use the following notation for the category of equivariant objects over $\mathcal{I}$ in $\mathcal{C}$ by
 […]
 3. Equivariant characteristic morphism of Injectives
-We study the characteristic morphism introduced in the previous chapter for the subcategory of injectives in a category with enough injectives.
-We start by discussing how we can reinterpret the characteristic morphism by the work of [Toe07].
-We then use this insight to prove that we can treat the full subcategory of injectives similarly to a tilting object.
+We study the characteristic morphism introduced in the previous chapter for the subcategory of injectives.
+We start by discussing how we can reinterpret Hochschild cohomology and the characteristic morphism by the work of [Toe07].
+We then use these insights to prove that we can treat the full subcategory of injectives similarly to a tilting object.
 3.1. Interpreting characteristic morphisms of a subcategory as pullback
-We use observations from [Toe07] to study in more depth the characteristic morphism
+We use observations from [Toe07] to study the characteristic morphism
 $c_{G}:\mathrm{HH}^{*}\left(\mathcal{C},\mathcal{M}\right)\to\mathrm{Ext}_{\mathcal{C}}^{*}\left(G,\mathcal{M}\otimes_{\mathcal{C}}G\right)$
 and the equivariant characteristic morphism
-$c_{\Gamma,G}:\mathrm{HH}^{*}\left(\mathcal{C},\mathcal{M}\right)\to\mathrm{Ext}_{\mathcal{C}_{\Gamma}}^{*}\left(G,\mathcal{M}\otimes_{\mathcal{C}}G\right).$
+$c_{\Gamma,T}:\mathrm{HH}^{*}\left(\mathcal{C},\mathcal{M}\right)\to\mathrm{Ext}_{\mathcal{C}_{\Gamma}}^{*}\left(T,\mathcal{M}\otimes_{\mathcal{C}}T\right).$
 We start by recalling from the previous Section 2.2 that an object $G\in\mathcal{C}$ is equivalently a functor from the one object category $𝕜*_{\Bbbk}$ with endomorphisms $𝕜\Bbbk$ into $\mathcal{C}$.
@@
 Definition 3.1.
-Let $\mathcal{C}$ be an abelian category and $I\subset\mathcal{C}$ a subcategory.
-Then denote by $\iota_{I}:I\hookrightarrow\mathcal{C}\subset\mathcal{D}_{\mathrm{dg}}^{b}\left(\mathcal{C}\right)$ the inclusion of objects of $I$.
+Let $\mathcal{C}$ be an abelian category and $\mathcal{I}\subset\mathcal{C}$ a subcategory.
+Then denote by $\iota_{I}:\mathcal{I}\hookrightarrow\mathcal{C}\subset\mathcal{D}_{\mathrm{dg}}^{b}\left(\mathcal{C}\right)$ the inclusion of $\mathcal{I}$.
 Remark 3.2.
@@
 In this interpretation, for a $\Gamma$-equivariant object $T$ the equivariant characteristic morphism turns into application of the pullback functor along the inclusion of the one object category $\iota_{T}:\Gamma\to\mathcal{D}_{\mathrm{dg}}^{b}\left(\mathrm{Qcoh}X\right)$:
-$\mathrm{HH}^{*}\left(\mathcal{C},\mathcal{M}\right)\cong\mathrm{Ho}\left(\mathrm{dg}-\mathrm{Cat}\right)\left(\mathcal{D}^{b}_{\mathrm{dg}}\left(\mathcal{C}\right),\mathcal{D}^{b}_{\mathrm{dg}}\left(\mathcal{C}\right)\right)\left(\mathrm{Id},\psi_{\mathcal{M}}\right)$$\mathrm{Ho}\left(\mathrm{dg}-\mathrm{Cat}\right)\left(\Gamma,\mathcal{D}^{b}_{\mathrm{dg}}\left(\mathcal{C}\right)\right)\left(\iota_{\mathcal{G}},\psi_{\mathcal{M}}\circ\iota_{\mathcal{G}}\right)\cong\mathrm{Ext}_{\mathcal{C}_{\Gamma}}\left(\mathcal{G},\mathcal{G}\otimes\mathcal{M}\right).$$c_{\Gamma,\mathcal{G}}\cong\iota_{\mathcal{G},*}$
-In particular we can use this interpretation to compute the equivariant characteristic morphism of the subcategory $I\in\mathcal{C}$:
-$\mathrm{HH}^{*}\left(\mathcal{C},\mathcal{M}\right)\cong\mathrm{Ho}\left(\mathrm{dg}-\mathrm{Cat}\right)\left(\mathcal{D}^{b}_{\mathrm{dg}}\left(\mathcal{C}\right),\mathcal{D}^{b}_{\mathrm{dg}}\left(\mathcal{C}\right)\right)\left(\mathrm{Id},\psi_{\mathcal{M}}\right)$$\mathrm{Ho}\left(\mathrm{dg}-\mathrm{Cat}\right)\left(\mathcal{I},\mathcal{D}^{b}_{\mathrm{dg}}\left(\mathcal{C}\right)\right)\left(\iota,\psi_{M}\circ\iota\right)\cong\mathrm{Ext}^{*}_{\mathcal{C}_{\mathcal{I}}}\left(\iota,\psi_{M}\circ\iota\right).$$c_{\iota_{I},I}=\iota_{I,*}$
+$\mathrm{HH}^{*}\left(\mathcal{C},\mathcal{M}\right)\cong\mathrm{Ho}\left(\mathrm{dg}\mathrm{Cat}\right)\left(\mathcal{D}^{b}_{\mathrm{dg}}\left(\mathcal{C}\right),\mathcal{D}^{b}_{\mathrm{dg}}\left(\mathcal{C}\right)\right)\left(\mathrm{Id},\psi_{\mathcal{M}}\right)$$\mathrm{Ho}\left(\mathrm{dg}\mathrm{Cat}\right)\left(\Gamma,\mathcal{D}^{b}_{\mathrm{dg}}\left(\mathcal{C}\right)\right)\left(\iota_{T},\psi_{\mathcal{M}}\circ\iota_{\mathcal{G}}\right)\cong\mathrm{Ext}_{\mathcal{C}_{\Gamma}}\left(T,T\otimes\mathcal{M}\right).$$c_{\Gamma,T}\cong\iota_{T,*}$
+In particular we can use this interpretation to compute the equivariant characteristic morphism of the subcategory $\mathcal{I}\subset\mathcal{C}$:
+$\mathrm{HH}^{*}\left(\mathcal{C},\mathcal{M}\right)\cong\mathrm{Ho}\left(\mathrm{dg}\mathrm{Cat}\right)\left(\mathcal{D}^{b}_{\mathrm{dg}}\left(\mathcal{C}\right),\mathcal{D}^{b}_{\mathrm{dg}}\left(\mathcal{C}\right)\right)\left(\mathrm{Id},\psi_{\mathcal{M}}\right)$$\mathrm{Ho}\left(\mathrm{dg}\mathrm{Cat}\right)\left(\mathcal{I},\mathcal{D}^{b}_{\mathrm{dg}}\left(\mathcal{C}\right)\right)\left(\iota,\psi_{M}\circ\iota\right)\cong\mathrm{Ext}^{*}_{\mathcal{C}_{\mathcal{I}}}\left(\iota,\psi_{M}\circ\iota\right).$$c_{\iota_{\mathcal{I}},\mathcal{I}}=\iota_{\mathcal{I},*}$
 The above observation is in particular useful as explicit computations for bimodules between $𝕜\Bbbk$-linear categories, or, even more, $\mathcal{A}_{\infty}$/$\mathrm{dg}$-categories can be very unwieldy.
-However, the above observation allows us to compute the equivariant characteristic morphisms of the $I$-$\mathcal{C}$ bimodule $\iota_{I}$ in special cases.
+However, the above observation allows us to compute the equivariant characteristic morphisms of the $\mathcal{I}$-$\mathcal{C}$ bimodule $\iota_{\mathcal{I}}$ in special cases.
 3.2. The $\mathrm{Inj}$-equivariant characteristic morphism
@@
 In particular we consider throughout this section the functor
-$\iota:\mathrm{Inj}\hookrightarrow\mathcal{C}\subset\mathcal{D}_{\mathrm{dg}}\left(\mathcal{C}\right)$
+$\iota:\mathrm{Inj}\hookrightarrow\mathcal{C}\subset\mathcal{D}_{\mathrm{dg}}\left(\mathcal{C}\right),$
+for $\mathcal{C}$ a category with enough injectives.
 Lemma 3.3.
-Let $\mathcal{C}$ be a $𝕜\Bbbk$-linear category with enough injectives. Then the inclusion $\iota:\mathrm{Inj}\hookrightarrow\mathcal{C}\subset\mathcal{D}_{\mathrm{dg}}\left(\mathcal{C}\right)$ is dg.
+Let $\mathcal{C}$ be a $𝕜\Bbbk$-linear category with enough injectives and finite global dimension.
+Then the inclusion $\iota_{*}:\mathrm{Inj}\hookrightarrow\mathcal{C}\subset\mathcal{D}^{b}_{\mathrm{dg}}\left(\mathcal{C}\right)$ is dg.
 Proof.
@@
 Proposition 3.4.
-Let $\mathcal{C}$ be a $𝕜\Bbbk$-linear category with enough injectives.
-Then the functor $\iota^{*}:\mathcal{D}_{\mathrm{dg}}\left(\mathcal{C}\right)\to\mathcal{D}_{\mathrm{dg}}\left(\mathrm{Inj}\right)$ is a quasi-equivalence.
+Let $\mathcal{C}$ be a $𝕜\Bbbk$-linear category with enough injectives and finite global dimension.
+Then the functor $\iota^{*}:\mathcal{D}^{b}_{\mathrm{dg}}\left(\mathcal{C}\right)\to\mathcal{D}^{b}_{\mathrm{dg}}\left(\mathrm{Inj}\right)$ is a quasi-equivalence.
 Proof.
@@
 Corollary 3.5.
-Let $\mathcal{C}$ be a $𝕜\Bbbk$-linear category with enough injectives and finite global dimension. Then the equivariant characteristic morphism $c_{\iota,\mathrm{Inj}}$ is an isomorphism.
+Let $\mathcal{C}$ be a $𝕜\Bbbk$-linear category with enough injectives and finite global dimension. Then the equivariant characteristic morphism
+$c_{\iota,\mathrm{Inj}}\cong\iota_{*}:\mathrm{HH}\left(\mathcal{C},\mathcal{M}\right)\to\mathrm{Ext}_{\mathcal{C}_{\mathrm{Inj}}}\left(\iota,\psi_{\mathcal{M}}\circ\iota\right)$
+is an isomorphism.
 Proof.
-By Section 3.1 above we can interpret the equivariant characteristic morphism as the pullback along the inclusion of the full subcategory $\iota:\mathrm{Inj}\to\mathcal{D}^{b}_{\mathrm{dg}}\left(\mathcal{C}\right)$ in the homotopy category of $\mathrm{dg}$-categories.
+By Section 3.1 above we can interpret the equivariant characteristic morphism as the pullback along the inclusion of the full subcategory
+$\iota:\mathrm{Inj}\to\mathcal{D}^{b}_{\mathrm{dg}}\left(\mathcal{C}\right)$
+in the homotopy category of $\mathrm{dg}$-categories.
 In particular we need to pick a fibrant replacement of $\mathrm{Inj}$.
 By [Tab07, Lemme 2.13] pretriangulated replacements are an incarnation of a fibrant replacement.
-So we may choose the canonical pretriangulated hull given by $\mathrm{dg}$-modules $\mathcal{D}_{\mathrm{dg}}\left(\mathrm{Inj}\right)$.
+So we may choose the canonical pretriangulated hull given by $\mathrm{dg}$-modules $\mathcal{D}^{b}_{\mathrm{dg}}\left(\mathrm{Inj}\right)$.
 By Proposition 3.4 $\iota$ induces an equivalence
-$\iota_{*}:\mathcal{D}_{\mathrm{dg}}\left(\mathcal{C}\right)\to\mathcal{D}_{\mathrm{dg}}\left(\mathrm{Inj}\right)$
-and hence in particular $c_{\iota,\mathrm{Inj}}\cong\iota_{*}$ is an isomorphism. ∎
+$\iota_{*}:\mathcal{D}^{b}_{\mathrm{dg}}\left(\mathcal{C}\right)\to\mathcal{D}^{b}_{\mathrm{dg}}\left(\mathrm{Inj}\right)$
+and hence in particular
+$c_{\iota,\mathrm{Inj}}\cong\iota_{*}:\mathrm{HH}^{*}\left(\mathcal{C},\mathcal{M}\right)\to\mathrm{Ext}^{*}_{\mathcal{C}_{\mathrm{Inj}}}\left(\iota,\psi_{\mathcal{M}}\circ\iota\right)$
+is an isomorphism. ∎
 Corollary 3.6.
@@
 We have that $\mathrm{Qcoh}\left(X\right)$ has enough injectives. In particular by Corollary 3.5 we have that
-$c_{\iota,\mathrm{Inj}}:\mathrm{HH}^{*}\left(\mathrm{Qcoh}\left(X\right),\mathcal{M}\right)\xrightarrow{\sim}\mathrm{Ho}\left(\mathrm{dg}-\mathrm{Cat}\right)\left(\mathrm{Inj},\mathcal{D}^{b}_{\mathrm{dg}}\left(\mathrm{Qcoh}\left(X\right)\right)\right)\left(\iota,\psi_{\mathcal{M}}\circ\iota\right)$
+$c_{\iota,\mathrm{Inj}}:\mathrm{HH}^{*}\left(\mathrm{Qcoh}\left(X\right),\mathcal{M}\right)\xrightarrow{\sim}\mathrm{Ho}\left(\mathrm{dg}\mathrm{Cat}\right)\left(\mathrm{Inj},\mathcal{D}^{b}_{\mathrm{dg}}\left(\mathrm{Qcoh}\left(X\right)\right)\right)\left(\iota,\psi_{\mathcal{M}}\circ\iota\right)$
 is an isomorphism. ∎
@@
 The above discussion shows that for smooth projective schemes the full subcategory of injectives behaves like a tilting object.
-Corollary 3.8.
-Let $\mathcal{C}$ be a $𝕜\Bbbk$-linear category with enough injectives, $\mathcal{M}$ a $𝕜\Bbbk$-central $\mathcal{X}$-bimodule and let $0\neq\eta\in\mathrm{HH}^{*}\left(\mathcal{C},\mathcal{M}\right)$.
-Then we have that its image under the $\mathrm{Inj}$-equivariant characteristic morphism $c_{\iota,\mathrm{Inj}}\left(\eta\right)$ does not vanish.
 4. Construction of $\psi_{\eta}$
 We recall the construction of the candidate non-Fourier-Mukai functor $\psi_{\eta}$.
-In particular from now on we will exclusively consider a functor of the shape $\psi_{\eta}$ for a smooth divisor $X\hookrightarrow Y$ with $\dim X=n$ and $0\neq\eta\in\mathrm{HH}^{m}\left(X,M\right)$ with $M$ a line-bundle, $m\geq n+3$ and $f_{*}\eta=0$.
 In order to apply Definition 4.1, Lemma 4.2 and Lemma 4.5 we assume that every quasi-projective scheme $X$ comes equipped with an open affine cover
@@
 [LVdB11, Definition 4.2] Let $X=\bigcup_{i=0}^{k}U_{i}$ be an open affine cover of a quasiprojective scheme $X$.
-Consider for $I\subset\left\{1,...,k\right\}$ the open sets $U_{I}:=\bigcap_{i\in I}U_{i}$.
+Consider for $\emptyset\neq I\subset\left\{1,...,k\right\}$ the open sets $U_{I}:=\bigcap_{i\in I}U_{i}$.
 Then $\mathcal{X}$ is the category with objects $\mathrm{P}\left\{1,...,k\right\}\setminus\emptyset$, morphisms:
 $\mathcal{X}\left(I,J\right):=\begin{cases}\mathcal{O}_{X}\left(U_{J}\right)&I\subset J\\ 0&\text{else}\end{cases}$
-and composition of morphisms induced by multiplication in the local ring composed with the restriction morphism.
+and composition of morphisms induced by multiplication in the local rings composed with the restriction morphism.
 One can think about $\mathcal{X}\mathrm{-mod}$ roughly as the category of presheaves associated to an affine open cover.
@@
 Lemma 4.2.
-[Kü22, Lemma 5.2] Let $X$ be smooth quasi-projective. Then there is a fully faithful embedding
+[Kü24, Lemma 5.2] Let $X$ be smooth quasi-projective. Then there is a fully faithful embedding
 $w:\mathcal{D}^{b}\left(\mathrm{Qcoh}X\right)\hookrightarrow\mathcal{D}^{b}\left(\mathcal{X}\right)$
@@
 Observe in particular that we have $W\mathcal{O}_{\Delta}\cong\mathcal{X}$ giving rise to the last equality in Lemma 4.2.
+Corollary 4.4.
+Let $\mathcal{X}$ be a $𝕜\Bbbk$-linear category associated to a scheme, let $\mathcal{M}=WM$ for $M$ a quasicoherent sheaf and let $0\neq\eta\in\mathrm{HH}^{*}\left(\mathcal{C},\mathcal{M}\right)$.
+Then we have that its image under the $\mathrm{Inj}$-equivariant characteristic morphism $c_{\iota,\mathrm{Inj}}\left(\eta\right)$ does not vanish.
 Lemma 4.5.
-[Kü22, Lemma 5.4] Let $f:X\hookrightarrow Y$ be a closed embedding of quasi projective schemes.
+[Kü24, Lemma 5.4] Let $f:X\hookrightarrow Y$ be a closed embedding of quasi projective schemes.
 Then we may pull back the open affine cover of $Y$ in the construction of $\mathcal{Y}$ and get an induced functor $\mathfrak{f}:\mathcal{Y}\to\mathcal{X}$ such that the diagram
@@
 As we will mostly be working with the category $\mathcal{D}^{b}\left(\mathcal{X}\right)$ and $w$ is an exact embedding we will denote the functor $\mathfrak{f}:\mathcal{X}\to\mathcal{Y}$ by $f:\mathcal{X}\to\mathcal{Y}$ for ease of readability.
-We also use the slightly strange convention of $f_{*}$ for the pullback on the $𝕜\Bbbk$-linear side in order to mimic the geometric setting and improve readability.
 The following is an application of [RVdBN19, Proposition 5.3.1] to our setting.
 Proposition 4.6.
-[Kü22, Proposition 5.5] Let $X$ be smooth projective of dimension $n$ and let $\eta\in\mathrm{HH}^{\geq n+3}\left(X,M\right)$. Then there exists an exact functor
+[Kü24, Proposition 5.5] Let $X$ be smooth projective of dimension $n$ and let $\eta\in\mathrm{HH}^{\geq n+3}\left(X,M\right)$. Then there exists an exact functor
 $L:\mathcal{D}^{b}\left(\mathrm{Qcoh}X\right)\to\mathcal{D}^{b}_{wQcoh\left(X\right)}\left(\mathcal{X}_{\eta}\right).$
@@
 […]
-Construction 4.7.
+Construction 4.8.
 [RVdBN19] Let $X\hookrightarrow Y$ be the inclusion of a smooth divisor into a smooth projective scheme and let
@@
 Remark 4.9.
-Construction 4.6 works out as well if $\eta$ is zero, however, the arising functor then trivially admits a lift to the $\mathcal{A}_{\infty}$-level and so we will not consider this case.
+Construction 4.7 works as well if $\eta$ is zero, however, the arising functor then trivially admits a lift to the $\mathcal{A}_{\infty}$-level and so we will not consider this case.
 5. Recollection on Obstructions
@@
 Definition 5.1.
-Let $\mathcal{X}$ be a small $𝕜\Bbbk$-linear category, assume $\mathcal{M}$ is right-flat, $\eta\in\mathrm{HH}^{\geq 3}\left(\mathcal{X},\mathcal{M}\right)$ and let $U\in\mathcal{X}\mathrm{-mod}$.
+Let $\mathcal{X}$ be a small $𝕜\Bbbk$-linear category, assume that $\mathcal{M}$ is a right-flat $𝕜\Bbbk$-central bimodule, $\eta\in\mathrm{HH}^{\geq 3}\left(\mathcal{X},\mathcal{M}\right)$ and let $U\in\mathcal{X}\mathrm{-mod}$.
 A lift of $U$ to $\mathcal{X}_{\eta}$ is an object $V\in\mathcal{D}^{b}_{\mathrm{dg}}\left(\mathcal{X}_{\eta}\right)$, such that there exists an isomorphism in $\mathcal{X}_{\eta}\mathrm{-mod}$:
 $\phi:\mathrm{H}^{*}\left(V\right)\xrightarrow{\sim}\mathrm{H}^{*}\left(\mathcal{X}_{\eta}\right)\otimes_{\mathcal{X}}U.$
-Dually, if $\mathcal{M}$ is left-projective and $U\in\mathcal{X}\mathrm{-mod}$.
+Dually, if $\mathcal{M}$ is a left-projective $𝕜\Bbbk$-central bimodule and $U\in\mathcal{X}\mathrm{-mod}$.
 A colift of $U$ to $\mathcal{X}_{\eta}$ is an object $V\in\mathcal{D}^{b}_{\mathrm{dg}}\left(\mathcal{X}_{\eta}\right)$, such that there exists an isomorphism in $\mathcal{X}_{\eta}\mathrm{-mod}$:
@@
 Lemma 5.5.
-Let $\iota\in\mathcal{D}\left(\mathcal{X}\otimes\mathcal{I}^{op}\right)$ for a small $𝕜\Bbbk$-linear category $\mathcal{I}$.
-Let $\widehat{\iota}\in\mathcal{D}_{\mathrm{dg}}\left(\mathcal{X}_{\eta}\otimes\mathcal{I}^{op}\right)$.
+Let $\iota\in\mathcal{D}^{b}\left(\mathcal{X}\otimes\mathcal{I}^{op}\right)$ for a small $𝕜\Bbbk$-linear category $\mathcal{I}$.
+Let $\widehat{\iota}\in\mathcal{D}^{b}_{\mathrm{dg}}\left(\mathcal{X}_{\eta}\otimes\mathcal{I}^{op}\right)$.
 Then the following are equivalent:
-  • The object $\widehat{\iota}$ is a colift of $\iota\in\mathcal{D}\left(\mathcal{X}\otimes\mathcal{I}^{op}\right)$.
-  • The object $\widehat{\iota}$ is an $\mathcal{A}_{\infty}$-lift of $L\circ\iota\in\mathcal{D}\left(\mathcal{X}_{\eta}\right)_{\mathcal{I}}$ i.e. $L\circ\iota:\mathcal{I}\to\mathcal{X}_{\eta}$.
+  • The object $\widehat{\iota}$ is a colift of $\iota\in\mathcal{D}^{b}\left(\mathcal{X}\otimes\mathcal{I}^{op}\right)$.
+  • The object $\widehat{\iota}$ is an $\mathcal{A}_{\infty}$-lift of $L\circ\iota\in\mathcal{D}^{b}\left(\mathcal{X}_{\eta}\right)_{\mathcal{I}}$ i.e. $L\circ\iota:\mathcal{I}\to\mathcal{X}_{\eta}$.
 Proof.
@@
 For this consider the following concatenation of natural isomorphisms:
-$\mathrm{Ext}^{*}_{\mathcal{X}\otimes\mathcal{I}^{op}}\left(\left(\mathcal{X}\otimes\mathcal{I}^{op}\right)_{\eta\cup 1},\iota\right)$ $\cong\mathrm{Ext}^{*}_{\mathcal{X}\otimes\mathcal{I}^{op}}\left(\mathcal{X}_{\eta}\otimes\mathcal{I}^{op},\iota\right)$
+$\mathrm{Ext}^{*}_{\mathcal{X}\otimes\mathcal{I}^{op}}\left(\left(\mathcal{X}\otimes\mathcal{I}^{op}\right)_{\eta\cup 1},\iota\right)$
+$\cong\mathrm{Ext}^{*}_{\mathcal{X}\otimes\mathcal{I}^{op}}\left(\mathcal{X}_{\eta}\otimes\mathcal{I}^{op},\iota\right)$
 $\cong\mathrm{Ext}^{*}_{\mathcal{X}}\left(\mathcal{X}_{\eta},\iota\right)$
@@
 By Corollary 4.4 we have that $c_{\iota,\mathrm{Inj}}\left(\eta\right)\neq 0.$
-In particular such a colift cannot exist, and hence no $\mathcal{A}_{\infty}$-lift as well. ∎
+In particular such a colift cannot exist, and hence there also cannot exist an $\mathcal{A}_{\infty}$-lift. ∎
 6. Structure of the functor $\widetilde{f}$
 It will be key for our conclusion that we carefully pick the space we consider for computing our obstructions, respectively understand the structure of the functor $\widetilde{f}$.
-In particular we will from now on fix the inclusion of a smooth divisor $f:X\hookrightarrow Y$, a pulled back linebundle $M=f^{*}\widehat{M}$ and $0\neq\eta\in\mathrm{HH}^{m}\left(X,M\right)$ with $m\geq\dim X+3$.
+In particular we will from now on fix the inclusion of a smooth divisor $f:X\hookrightarrow Y$, a pulled back linebundle $M=f^{*}\widehat{M}$ and $0\neq\eta\in\mathrm{HH}^{m}\left(X,M\right)$ with $m\geq\dim X+3$ such that $f_{*}\eta$.
 We start of by computing the images of sheaves $\mathcal{F}\in\mathrm{Qcoh}\left(X\right)$ under the functor $\psi_{\eta}$:
@@
 Proof.
-Consider the distinguished triangle in $\mathcal{D}^{b}\left(\mathcal{X}_{\eta}\right)$ from [RVdBN19, Lemma 11.3]
+Consider the distinguished triangle in $\mathcal{D}^{b}\left(\mathcal{X}_{\eta}\right)$ induced by $L$, see [RVdBN19, Lemma 11.3]
 $\mathrm{can}_{*}\mathcal{F}\to L\mathcal{F}\to\mathrm{can}_{*}\Sigma^{2-m}\left(\mathcal{F}\otimes\mathcal{M}\right)\to\Sigma\mathrm{can}_{*}\mathcal{F},$
@@
 Now, as $\mathcal{F}$ and $\mathcal{M}$ are sheaves we have
-$\mathrm{Ext}^{1}_{\mathcal{Y}}\left(\Sigma^{2-m}f_{*}\mathcal{F}\otimes\mathcal{M},f_{*}\mathcal{F}\right)$ $\cong\mathrm{Ext}^{m-1}_{\mathcal{Y}}\left(f_{*}\mathcal{F}\otimes\mathcal{M},f_{*}\mathcal{F}\right)$ $\mathrm{Ext}^{i}\left(\Sigma^{-j}\_,\_\right)\cong\mathrm{Ext}^{i+j}\left(\_,\_\right)$
+$\mathrm{Ext}^{1}_{\mathcal{Y}}\left(\Sigma^{2-m}f_{*}\mathcal{F}\otimes\mathcal{M},f_{*}\mathcal{F}\right)$
+$\cong\mathrm{Ext}^{m-1}_{\mathcal{Y}}\left(f_{*}\mathcal{F}\otimes\mathcal{M},f_{*}\mathcal{F}\right)$
 $\cong\mathrm{Ext}^{m-1}_{Y}\left(f_{*}\mathcal{F}\otimes\mathcal{M},f_{*}\mathcal{F}\right)$ Lemma 4.2
@@
 In particular we can compute
-$\mathrm{Ext}_{Y}^{*}\left(\psi_{\eta}I,\psi_{\eta}J\right)$ $\cong\mathrm{Ext}_{Y}^{*}\left(f_{*}I\oplus\left(f_{*}I\otimes f_{*}M\left[2-m\right]\right),f_{*}J\oplus f_{*}\left(J\otimes M\left[2-m\right]\right)\right)$
-$\cong\begin{pmatrix}\mathrm{Ext}_{Y}^{*}\left(f_{*}I,f_{*}J\right)&\mathrm{Ext}_{Y}^{*}\left(f_{*}I,f_{*}\left(J\otimes M\right)[2-m]\right)\\ \mathrm{Ext}_{Y}^{*}\left(f_{*}\left(I\otimes M\right)[2-m],f_{*}J\right)&\mathrm{Ext}_{Y}^{*}\left(f_{*}I,f_{*}J\right)\end{pmatrix}$
+$\mathrm{Ext}_{Y}^{*}\left(\psi_{\eta}I,\psi_{\eta}J\right)$ $\cong\mathrm{Ext}_{Y}^{*}\left(f_{*}I\oplus\left(f_{*}I\otimes\widehat{M}\right)\left[2-m\right],f_{*}J\oplus\left(f_{*}J\otimes\widehat{M}\right)\left[2-m\right]\right)$
+$\cong\begin{pmatrix}\mathrm{Ext}_{Y}^{*}\left(f_{*}I,f_{*}J\right)&\mathrm{Ext}_{Y}^{*}\left(f_{*}I,f_{*}\left(J\right)\otimes\widehat{M}[2-m]\right)\\ \mathrm{Ext}_{Y}^{*}\left(f_{*}\left(I\right)\otimes\widehat{M}[2-m],f_{*}J\right)&\mathrm{Ext}_{Y}^{*}\left(f_{*}I,f_{*}J\right)\end{pmatrix}$
 Where we applied in the bottom right corner the fact that $\widehat{M}$ is invertible while also dropping the shifts. Now it suffices to observe that we have
-$Ext_{Y}^{*}\left(f_{*}F,f_{*}G\right)$ $\cong\mathrm{Ext}^{*}_{X}\left(f^{*}f_{*}F,G\right)$
-$\cong\mathrm{Ext}^{*}_{X}\left(I\otimes C_{f}[-1]\oplus F,G\right)$ $f^{*}f_{*}F\cong\left(F\oplus F\otimes C_{f}[1]\right)$
-$\cong\mathrm{Ext}^{*}_{X}\left(F,G\right)\oplus\mathrm{Ext}^{*}_{X}\left(F\otimes C_{f},G\right)$ $\mathrm{Ext}^{*}\text{ is additive}$
+$\mathrm{Ext}_{Y}^{*}\left(f_{*}F,f_{*}G\right)$ $\cong\mathrm{Ext}^{*}_{X}\left(f^{*}f_{*}F,G\right)$
+$\cong\mathrm{Ext}^{*}_{X}\left(I\otimes C_{f}[-1]\oplus F,G\right)$ $\mathllap{f^{*}f_{*}F\cong\left(F\oplus F\otimes C_{f}[1]\right)}$
+$\cong\mathrm{Ext}^{*}_{X}\left(F,G\right)\oplus\mathrm{Ext}^{*}_{X}\left(F\otimes C_{f},G\right)$ $\quad\quad\mathrm{Ext}^{*}\text{ is additive}$
 Where we use $C_{f}$ for the conormal bundle of the inclusion $f:X\hookrightarrow Y$.
@@
 ∎
-In order to have control over the obstructions explicitly, we will construct the functor $\widetilde{f}:\mathcal{Y}\to\mathcal{X}_{\eta}$ from Lemma 4.6 explicitly.
-For this consider the commutative diagram of $\mathcal{A}_{\infty}$-categories and functors giving rise to a good candidate.
+In order to have control over the obstructions explicitly, we will study the functor $\widetilde{f}:\mathcal{Y}\to\mathcal{X}_{\eta}$ from Lemma 4.7 explicitly.
+For this consider the commutative diagram of $\mathcal{A}_{\infty}$-categories and functors giving rise to a candidate.
 […]
-The bottom arrow is the functor given in Lemma 4.4 and we denoted by can the canonical functors arising from the infinitesimal deformation.
+The bottom arrow is the functor given in Lemma 4.5 and we denote by can the canonical functors arising from the infinitesimal deformation, see Definition 2.5.
 The other functors are given by:
   • The functor $f:\mathcal{Y}_{f_{*}\eta}\to\mathcal{X}_{\eta}$ is the functor given by the identity on objects and acting on morphisms by
-    $\begin{pmatrix}f&0\\ 0&f\end{pmatrix}:\mathcal{Y}_{f_{*}\eta}\left(i,j\right)=\mathcal{Y}\left(i,j\right)\oplus f_{*}\mathcal{M}\left(i,j\right)[2-m]\to\mathcal{X}\left(i,j\right)\oplus\mathcal{M}\left(i,j\right)[2-m].$
+    $\begin{pmatrix}f&0\\ 0&f\end{pmatrix}:\mathcal{Y}\left(i,j\right)\oplus f_{*}\mathcal{M}\left(i,j\right)[2-m]\to\mathcal{X}\left(i,j\right)\oplus\mathcal{M}\left(i,j\right)[2-m].$
     This functor is by construction an $\mathcal{A}_{\infty}$-functor as by definition $f_{*}\eta$ is given by the composition of bimodule morphisms
@@
 […]
-We then use the explicit structure of the category $f_{*}\mathrm{Inj}_{f_{*}\eta}$ to compute that these obstructions cannot vanish independent of intermediate lifts.
-Conceptually one can think about this argument as the bimodule $\psi_{\eta}$ not being $\mathcal{A}_{\infty}$ over $\mathrm{Inj}$, however, it is $\mathcal{A}_{\infty}$ over $\mathcal{Y}$.
-In particular it behaves like a non-$\mathcal{A}_{\infty}$ Fourier-Mukai kernel
-The following Propositions 6.5 and Proposition 6.6 shows that the notations $\mathrm{Inj}_{\eta}$ and $f_{*}\mathrm{Inj}_{f_{*}\eta}$ for the essential image of $f_{*}\circ L\circ\iota$ are warranted.
+Conceptually one can think about this argument as the bimodule $\psi_{\eta}$ not being $\mathcal{A}_{\infty}$ over $\mathrm{Inj}$ while being $\mathcal{A}_{\infty}$ over $\mathcal{Y}$.
+In particular it behaves like a non-$\mathcal{A}_{\infty}$ Fourier-Mukai kernel.
+The following Proposition 6.5 and Proposition 6.6 show that the notations $\mathrm{Inj}_{\eta}$ and $f_{*}\mathrm{Inj}_{f_{*}\eta}$ for the essential image of $f_{*}\circ L\circ\iota$ are warranted.
 Proposition 6.5.
@@
 Proof.
-By construction of $L$, in particular [RVdBN19, (5.7)], the morphism spaces in $\mathrm{Inj}_{\eta}$ are of the shape
+By construction of $L$, see [RVdBN19, (5.7)], the morphism spaces in $\mathrm{Inj}_{\eta}$ are of the shape
 $\mathrm{essim}\left(L\circ\iota\right)\left(I,J\right)$ $\cong\mathrm{Ext}_{X}^{*}\left(I,J\right)\oplus\mathrm{Ext}_{X}^{2-m+*}\left(I,J\otimes\mathcal{M}\right)$
@@
 Proposition 6.6.
-The essential image of $f_{*}\circ L\circ\iota$ has objects $i\in I$ and morphism spaces given by
+The essential image of $f_{*}\circ L\circ\iota$ has objects $I\in\mathrm{Inj}$ and morphism spaces given by
 $\textrm{essim}\left(f_{*}\circ L\circ\iota\right)\left(I,J\right)\cong\mathrm{Ext}_{Y}^{*}\left(f_{*}I,f_{*}J\right)\oplus\mathrm{Ext}_{Y}^{2-m+*}\left(f_{*}I,f_{*}\left(J\otimes\mathcal{M}\right)\right).$
@@
 By Lemma 6.1 we have $\psi_{\eta}\left(I\right)\cong f_{*}I\oplus f_{*}\left(M\otimes I\right)[2-m]$.
-As we made the choice for $s$ being the trivial inclusion $\mathcal{Y}\hookrightarrow\mathcal{Y}_{0}\cong\mathcal{Y}\oplus f_{*}\mathcal{M}[2-m]$, $s_{*}$ induces the inclusions of upper triangular matrices with identical diagonal, as claimed.
+As we made the choice for $s$ being the trivial inclusion $\mathcal{Y}\hookrightarrow\mathcal{Y}_{0}\cong\mathcal{Y}\oplus f_{*}\mathcal{M}[2-m]$, $s_{*}$ induces the inclusions of upper triangular matrices with identical diagonal as claimed.
 ∎
@@
 7. Computing obstructions explicitly for Divisors
-The observations from the previous section let us compute:
+The observations from the previous sections let us compute:
 Proposition 7.1.
@@
 We know by Proposition 5.6 that $L\circ\iota$ does not have an $\mathcal{A}_{\infty}$-lift. In particular we have by Lemma 5.4 that there must be a non-vanishing
-$o_{i}\left(L\circ\iota\right)\in\mathrm{HH}^{i}\left(\mathrm{Inj},\mathrm{Inj}_{\eta}\right)_{i}$
+$o_{i}\left(L\circ\iota\right)\in\mathrm{HH}^{i}\left(\mathrm{Inj},\mathrm{Inj}_{\eta}\right)_{i}.$
 However, as $L\circ\iota$ is a functor it is $\mathcal{A}_{2}$ and the non-vanishing obstruction must have $i\geq 3$.
@@
 $\varphi=\left(\varphi_{1},...,\varphi_{i}\right):\mathcal{I}\to f_{*}\mathrm{Inj}_{f_{*}\eta}$
-be an $\mathcal{A}_{i}$ functor from a $𝕜\Bbbk$-linear category $I$ then we have an isomorphism
+be an $\mathcal{A}_{i}$ functor from a $𝕜\Bbbk$-linear category $\mathcal{I}$ then we have an isomorphism
 $\left(s_{*}\circ h_{*}\circ j\right)_{*}:\mathrm{HH}^{*}\left(\mathcal{I},f_{*}\mathrm{Inj}_{f_{*}\eta}\right)_{2-m}$ $\xrightarrow{\sim}\mathrm{HH}^{*}\left(\mathcal{I},\mathcal{D}^{b}\left(\mathrm{Qcoh}Y\right)\right)$
-$o_{i}\left(f_{*}\circ\varphi\right)$ $\mapsto o_{i}\left(s_{*}\circ h_{*}\circ j\circ\varphi\circ\iota\right).$
+$o_{i}\left(f_{*}\circ\varphi\right)$ $\mapsto o_{i}\left(s_{*}\circ h_{*}\circ j\circ\varphi\right).$
 Furthermore an $\mathcal{A}_{i+1}$-lift $\varphi_{i+1}$ of $s_{*}\circ h_{*}\circ j\circ\varphi$ gives rise to an $\mathcal{A}_{i+1}$-functor $\widehat{\varphi_{i+1}}$ such that
@@
 As $h$ is a homotopy equivalence and $j$ is the inclusion of a full subcategory we have by naturality of obstructions
-$\left(h\circ j\right)_{*}:\mathrm{HH}^{m}\left(\mathrm{Inj},f_{*}\mathrm{Inj}_{f_{*}\eta}\right)$ $\xrightarrow{\sim}\mathrm{HH}^{m}\left(\mathrm{Inj},\mathcal{D}^{b}\left(\mathcal{Y}_{0}\right)\right)$
-$o_{m}\left(\varphi\right)$ $\mapsto o_{m}\left(h_{*}\circ j\circ\varphi\iota\right).$
-For the compatibility with lifts observe that since $j$ and $h$ are homotopy equivalences on the image of the injectives and so admit quasi-inverses they are immediately compatible with intermediate lifts.
+$\left(h\circ j\right)_{*}:\mathrm{HH}^{m}\left(\mathcal{I},f_{*}\mathrm{Inj}_{f_{*}\eta}\right)$ $\xrightarrow{\sim}\mathrm{HH}^{m}\left(\mathcal{I},\mathcal{D}^{b}\left(\mathcal{Y}_{0}\right)\right)$
+$o_{m}\left(\varphi\right)$ $\mapsto o_{m}\left(h_{*}\circ j\circ\varphi\right).$
+For the compatibility with lifts observe that since $j$ and $h$ are homotopy equivalences on the image of the injectives they admit quasi-inverses they are immediately compatible with intermediate lifts.
 As by Proposition 6.7 the functor $s_{*}$ is the inclusion of the upper triangular matrices and by Proposition 6.8 an equivalence on the negative $\mathrm{Ext}$-groups.
-We get, since $\mathrm{Inj}$ is a $𝕜\Bbbk$-linear category
-$s_{*}:\mathrm{HH}^{m}\left(\mathrm{Inj},\mathcal{D}^{b}\left(\mathcal{Y}_{0}\right)\right)_{<0}$ $\xrightarrow{\sim}\mathrm{HH}^{m}\left(\mathrm{Inj},\mathcal{D}^{b}\left(\mathrm{Qcoh}Y\right)\right)_{<0}$
+We get, since $\mathcal{I}$ is a $𝕜\Bbbk$-linear category that $s_{*}$ is an equivalence,
+$s_{*}:\mathrm{HH}^{m}\left(\mathcal{I},\mathcal{D}^{b}\left(\mathcal{Y}_{0}\right)\right)_{<0}$ $\xrightarrow{\sim}\mathrm{HH}^{m}\left(\mathcal{I},\mathcal{D}^{b}\left(\mathrm{Qcoh}Y\right)\right)_{<0}$
 $o_{m}\left(h_{*}\circ j\circ\varphi\right)$ $\mapsto o_{m}\left(s_{*}\circ h_{*}\circ j\circ\varphi\right).$
@@
 Lemma 7.6.
-Let $f^{*}\dashv f_{*}:\mathcal{D}_{\infty}\left(\mathrm{Qcoh}X\right)\to\mathcal{D}_{\infty}\left(\mathrm{Qcoh}Y\right)$ be an adjuction of $\mathcal{A}_{\infty}$-functors and let $N$ be an $\mathcal{A}_{\infty}$-$I$-bimodule of the shape
-$N:I\otimes I^{\mathrm{op}}$ $\to\mathcal{D}^{b}\left(\mathrm{Qcoh}X\right)$
+Let $f^{*}\dashv f_{*}:\mathcal{D}^{b}_{\infty}\left(\mathrm{Qcoh}X\right)\to\mathcal{D}_{\infty}^{b}\left(\mathrm{Qcoh}Y\right)$ be an adjuction of $\mathcal{A}_{\infty}$-functors and let $N$ be an $\mathcal{A}_{\infty}$$-\mathcal{I}$-bimodule of the shape
+$N:\mathcal{I}\otimes\mathcal{I}^{\mathrm{op}}$ $\to\mathcal{D}^{b}_{\infty}\left(\mathrm{Qcoh}X\right)$
 $\left(i,j\right)$ $\mapsto\mathrm{Ext}^{*}_{X}\left(Ni,Nj\right)$
 $\alpha\otimes\beta$ $\mapsto N_{1}\left(\alpha\right)\otimes N_{2}\left(\beta\right)$
-for $I$ a $𝕜\Bbbk$-linear category. Then we have that the adjunction quasi-isomorphism
+for $\mathcal{I}$ a $𝕜\Bbbk$-linear category. Then we have that the adjunction quasi-isomorphism
 $\psi:\mathrm{Ext}^{*}_{\mathcal{B}}\left(f_{*}X,f_{*}Y\right)\xrightarrow{\sim}\mathrm{Ext}^{*}_{\mathcal{A}}\left(f^{*}f_{*}X,Y\right)$
-induces a quasi-isomorphism of $\mathcal{A}_{\infty}$-$I$-bimodules.
-$\begin{pmatrix}f_{*}N:I\otimes I^{\mathrm{op}}\to\mathcal{D}^{b}\left(\mathrm{Qcoh}Y\right)\\ \left(i,j\right)\mapsto\mathrm{Ext}^{*}_{X}\left(f_{*}Ni,f_{*}Nj\right)\\ \alpha\otimes\beta\mapsto f_{*}N_{1}\left(\alpha\right)\otimes f_{*}N_{2}\left(\beta\right)\end{pmatrix}\cong\begin{pmatrix}I\otimes I^{\mathrm{op}}\to\mathcal{D}^{b}\left(\mathrm{Qcoh}X\right)\\ \left(i,j\right)\mapsto\mathrm{Ext}^{*}_{X}\left(f^{*}f_{*}Ni,Nj\right)\\ \alpha\otimes\beta\mapsto f^{*}f_{*}N_{1}\left(\alpha\right)\otimes N_{2}\left(\beta\right)\end{pmatrix}.$
+induces a quasi-isomorphism of $\mathcal{A}_{\infty}$$-\mathcal{I}$-bimodules.
+$\begin{pmatrix}f_{*}N:\mathcal{I}\otimes\mathcal{I}^{\mathrm{op}}\to\mathcal{D}^{b}\left(\mathrm{Qcoh}Y\right)\\ \left(i,j\right)\mapsto\mathrm{Ext}^{*}_{X}\left(f_{*}Ni,f_{*}Nj\right)\\ \alpha\otimes\beta\mapsto f_{*}N_{1}\left(\alpha\right)\otimes f_{*}N_{2}\left(\beta\right)\end{pmatrix}\cong\begin{pmatrix}\mathcal{I}\otimes\mathcal{I}^{\mathrm{op}}\to\mathcal{D}^{b}\left(\mathrm{Qcoh}X\right)\\ \left(i,j\right)\mapsto\mathrm{Ext}^{*}_{X}\left(f^{*}f_{*}Ni,Nj\right)\\ \alpha\otimes\beta\mapsto f^{*}f_{*}N_{1}\left(\alpha\right)\otimes N_{2}\left(\beta\right)\end{pmatrix}.$
 Proof.
 Consider a $\mathrm{dg}$-replacement of the adjunction
-$f^{*}\dashv f_{*}:\mathcal{D}_{\mathrm{dg}}\left(\mathrm{Qcoh}X\right)\to\mathcal{D}_{\mathrm{dg}}\left(\mathrm{Qcoh}Y\right).$
+$f^{*}\dashv f_{*}:\mathcal{D}^{b}_{\mathrm{dg}}\left(\mathrm{Qcoh}X\right)\to\mathcal{D}^{b}_{\mathrm{dg}}\left(\mathrm{Qcoh}Y\right).$
 Then it suffices to show that the isomorphism
-$\mathcal{D}_{\mathrm{dg}}\left(\mathrm{Qcoh}Y\right)\left(f_{*}Ni,f_{*}Nj\right)\xrightarrow{\sim}\mathcal{D}_{\mathrm{dg}}\left(\mathrm{Qcoh}X\right)\left(f^{*}f_{*}Ni,Nj\right)$
-is a quasi-isomorphism of $\mathrm{dg}$-$I$-bimodules.
+$\mathcal{D}^{b}_{\mathrm{dg}}\left(\mathrm{Qcoh}Y\right)\left(f_{*}Ni,f_{*}Nj\right)\xrightarrow{\sim}\mathcal{D}^{b}_{\mathrm{dg}}\left(\mathrm{Qcoh}X\right)\left(f^{*}f_{*}Ni,Nj\right)$
+is a quasi-isomorphism of $\mathrm{dg}$$-\mathcal{I}$-bimodules.
 By definition of $\mathrm{dg}$-adjunction it is already a quasi-isomorphism of chain complexes of vectorspaces.
-In particular it suffices to prove that it is a morphism of $I$-bimodules.
+In particular it suffices to prove that it is a morphism of $\mathcal{I}$-bimodules.
 For this consider first the operation from the left on $f_{*}N$. Denote it by
-$f_{*}N_{1}\left(\alpha\right):f_{*}Ni^{\prime}\to f_{*}Ni$
-, then we have the commutative diagram
+$f_{*}N_{1}\left(\alpha\right):f_{*}Ni^{\prime}\to f_{*}Ni,$
+then we have the commutative diagram
 […]
@@
 In particular
-$\mathcal{D}_{\mathrm{dg}}\left(\mathrm{Qcoh}\left(Y\right)\right)\left(f_{*}Ni,f_{*}Nj\right)\xrightarrow{\sim}\mathcal{D}_{\mathrm{dg}}\left(\mathrm{Qcoh}X\right)\left(f^{*}f_{*}Ni,Nj\right)$
-is compatible with both actions and induces a quasi-isomorphism of $\mathcal{A}_{\infty}$-$I$-bimodules. ∎
+$\mathcal{D}^{b}_{\mathrm{dg}}\left(\mathrm{Qcoh}\left(Y\right)\right)\left(f_{*}Ni,f_{*}Nj\right)\xrightarrow{\sim}\mathcal{D}^{b}_{\mathrm{dg}}\left(\mathrm{Qcoh}X\right)\left(f^{*}f_{*}Ni,Nj\right)$
+is compatible with both actions and induces a quasi-isomorphism of $\mathcal{A}_{\infty}$$-\mathcal{I}$-bimodules. ∎
 Remark 7.7.
@@
 Otherwise we would not be able to compare the right action, as not every functor in the essential image arises via $f_{*}$.
-The following is a rewritten proof to our setup of the well known fact that for a divisor $X\hookrightarrow Y$ the object $f_{*}\mathcal{F}\in\mathcal{D}_{\infty}\left(Y\right)$ usually is not formal over $Y$ for a sheaf $\mathcal{F}$, however, it is formal over $X$.
+The following is a rewritten proof to our setup of the well known fact that for a divisor $X\hookrightarrow Y$ and a sheaf $\mathcal{F}$ the object $f_{*}\mathcal{F}\in\mathcal{D}^{b}_{\mathrm{dg}}\left(Y\right)$ usually is not formal over $Y$, however, it is formal over $X$.
 In particular it is essential that we restrict in this proof to the bimodule structure over $\mathrm{Inj}$.
@@
 $\cong\mathrm{Ext}^{*-2+m}_{X}\left(f^{*}f_{*}\mathcal{O}_{X}\otimes_{\mathcal{O}_{Y}}I,J\otimes M\right)$
-$\cong\mathrm{Ext}^{*-2+m}_{X}\left(I\otimes C_{f}[1]\oplus I,J\otimes M\right)$ $f^{*}f_{*}\mathcal{O}_{X}\cong\otimes\left(\mathcal{O}_{X}\oplus C_{f}[1]\right)$
+$\cong\mathrm{Ext}^{*-2+m}_{X}\left(I\otimes C_{f}[1]\oplus I,J\otimes M\right)$ $\mathllap{f^{*}f_{*}\mathcal{O}_{X}\cong\otimes\left(\mathcal{O}_{X}\oplus C_{f}[1]\right)}$
 $\cong\mathrm{Ext}^{*-2+m}_{X}\left(I,J\otimes M\right)\oplus\mathrm{Ext}^{*-3+m}_{X}\left(I\otimes C_{f},J\otimes M\right)$ $\mathrm{Ext}^{*}\text{ is additive}$
@@
 $\mathrm{m}_{3}=0$
-on the $\mathrm{Inj}$-bimodule $\mathrm{H}^{<0}\mathrm{\mathrm{im}}f_{*}\circ L\circ\iota$.
-Now we compute in $\mathrm{HH}^{m}\left(\mathrm{Inj},\mathrm{H}^{2-m}\mathrm{\mathrm{im}}\psi_{\eta}\circ\iota\right)$ that the obstruction is independent of choices
+on the $\mathrm{Inj}$-bimodule $\mathrm{H}^{<0}\mathrm{\mathrm{im}}f_{*}\circ L\circ\iota$. Now we compute in
+$\mathrm{HH}^{m}\left(\mathrm{Inj},\mathrm{H}^{2-m}\mathrm{\mathrm{im}}\psi_{\eta}\circ\iota\right)=\mathrm{HH}^{m}\left(\mathrm{Inj},\mathrm{H}^{2-m}\mathrm{\mathrm{im}}^{<0}\psi_{\eta}\circ\iota\right)$
+that the obstruction is independent of choices
 […]
@@
 Where we used in the last line that $f_{*}$ induces an equivalence by Lemma 7.4 on
-$\mathrm{Ext}_{\mathcal{X}_{\eta}}^{2-m}\left(L\left(I\right),L\left(J\right)\right)\cong\mathrm{Ext}_{\mathcal{Y}_{f_{*}\eta}}^{2-m}\left(f_{*}\circ L\left(I\right),f_{*}\circ L\left(J\right)\right)$
+$\mathrm{Ext}_{\mathcal{X}_{\eta}}^{2-m}\left(L\left(I\right),L\left(J\right)\right)\cong\mathrm{Ext}_{\mathcal{Y}_{f_{*}\eta}}^{2-m}\left(f_{*}\circ L\left(I\right),f_{*}\circ L\left(J\right)\right).$
 In particular the obstruction does not vanish and is independent of choices. ∎
@@
 $o_{m}\left(\psi_{\eta}\circ\iota\right)=o_{m}\left(f_{*}\circ L\circ\iota\right)$
-and by Lemma 7.10
+and by Lemma 7.10 we have independent of choices
 $o_{m}\left(f_{*}\circ L\circ\iota\right)\neq 0.$
@@
 Theorem 7.12.
-The functor $\psi_{\eta}$ does not admit an $\mathcal{A}_{\infty}$/$\mathrm{dg}$-lift.
+The functor
+$\psi_{\eta}:\mathcal{D}^{b}\left(\mathrm{Qcoh}X\right)\to\mathcal{D}^{b}\left(\mathrm{Qcoh}Y\right)$
+does not admit an $\mathcal{A}_{\infty}$/$\mathrm{dg}$-lift.
 Proof.
@@
 In particular $\psi_{\eta}\circ\iota$ cannot have an $\mathcal{A}_{\infty}$-lift.
-Since $\iota$ is dg, this means that $\psi_{\eta}$ cannot have an $\mathcal{A}_{\infty}$-lift.
+Since $\iota$ is dg by Lemma 3.3, this means that $\psi_{\eta}$ cannot have an $\mathcal{A}_{\infty}$-lift.
 ∎
-By [Kü22, Theorem 4.14] we can compute for a smooth hypersurface the kernels of $\mathrm{HH}^{*}\left(X,\mathcal{O}\left(p\right)\right)$ explicitly via twisted Hodge-diamonds, in particular all these kernels in $\mathrm{HH}^{\geq\dim X+3}\left(X,\mathcal{O}\left(p\right)\right)$ define non-$\mathrm{dg}$ functors.
+If the functor $\psi_{\eta}$ is continuous we can restrict our argument to the category of coherent sheaves.
+We conjecture that $\psi_{\eta}$ has this property as it is by definition defined on the subcategory of injectives and extended from there, see [RVdBN19].
+However, as it is very ill-behaved we were not yet able to verify it.
+Theorem 7.13.
+Assume that $\psi_{\eta}$ is continuous. Then we have that the restriction
+$\psi_{\eta}:\mathcal{D}^{b}\left(X\right)\to\mathcal{D}^{b}\left(Y\right)$
+is not Fourier-Mukai.
+Proof.
+Assume that $\psi_{\eta}:\mathcal{D}^{b}\left(X\right)\to\mathcal{D}^{b}\left(Y\right)$ admits a dg-enhancement.
+Let $\widehat{\psi}_{\eta}:\mathcal{D}^{b}\left(\mathrm{Qcoh}X\right)\to\mathcal{D}^{b}\left(\mathrm{Qcoh}Y\right)$ be the continuation of $\varphi_{\eta}$ to $\mathcal{D}^{b}\left(\mathrm{Qcoh}\_\right)$ via [Lur09, Proposition 5.3.5.10.], respectively [Toe07, Theorem 7.2], we have that this is the unique continuous extension of $\psi_{\eta}$.
+In particular we have, as $\psi_{\eta}$ was assumed to be continuous
+$\widehat{\psi_{\eta}}=\psi_{\eta}::\mathcal{D}^{b}\left(\mathrm{Qcoh}X\right)\to\mathcal{D}^{b}\left(\mathrm{Qcoh}Y\right).$
+But as $\psi_{\eta}:\mathcal{D}^{b}\left(X\right)\to\mathcal{D}^{b}\left(Y\right)$ had a $\mathrm{dg}$-enhancement, $\widehat{\psi_{\eta}}$ would have a $\mathrm{dg}$-enhancement, which is contradiction to Theorem 7.12
+This in particular means that the following corollaries could be restricted to the case of coherent sheaves if $\psi_{\eta}$ is continuous.
+By [Kü24, Theorem 4.14] we can compute for a smooth hypersurface the kernels of $\mathrm{HH}^{*}\left(X,\mathcal{O}\left(p\right)\right)$ explicitly via twisted Hodge-diamonds, in particular all these kernels in $\mathrm{HH}^{\geq\dim X+3}\left(X,\mathcal{O}\left(p\right)\right)$ define non-$\mathrm{dg}$ functors.
 Corollary 7.14.
@@
 $\begin{matrix}\;&\;&\;&\;&\;&\;&\;&1&\;&\;&\;&\;&\;\\ \;&\;&\;&\;&\;&\;&0&\;&0&\;&\;&\;&\;\\ \;&\;&\;&\;&\;&0&\;&1&\;&0&\;&\;&\;\\ \;&\;&\;&\;&0&\;&0&\;&0&\;&0&\;&\;\\ \;&\;&\;&0&\;&0&\;&1&\;&0&\;&0&\;\\ \;&\;&0&\;&0&\;&0&\;&0&\;&0&\;&0\\ \;&0&\;&\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\mathclap{36}\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}&\;&\mathclap{2472}&\;&\mathclap{8093}&\;&\mathclap{2472}&\;&\mathclap{36}&\;&0\\ \;&\;&0&\;&0&\;&0&\;&0&\;&0&\;&0\\ \;&\;&\;&0&\;&0&\;&1&\;&0&\;&0&\;&\;&\;\\ \;&\;&\;&\;&0&\;&0&\;&0&\;&0&\;&\;&\;&\;\\ \;&\;&\;&\;&\;&0&\;&1&\;&0&\;&\;&\;&\;&\;\\ \;&\;&\;&\;&\;&\;&0&\;&0&\;&\;&\;&\;&\;&\;\\ \;&\;&\;&\;&\;&\;&\;&1.&\;&\;&\;&\;&\;&\;&\;\\ \end{matrix}$
-And so we have by [Kü22, Theorem 4.14],
+And so we have by [Kü24, Theorem 4.14],
 $\dim\ker\left(f_{*}:\mathrm{HH}^{10}\left(X,\mathcal{O}_{X}\left(1\right)\right)\to\mathrm{HH}^{10}\left(\mathbb{P}^{7},f_{*}\mathcal{O}_{X}\left(1\right)\right)\right)$ $=36$
@@
 $\begin{matrix}\;&\;&\;&\;&\;&0&\;&\;&\;&\;&\;\\ \;&\;&\;&\;&0&\;&0&\;&\;&\;&\;\\ \;&\;&\;&0&\;&0&\;&0&\;&\;&\;\\ \;&\;&0&\;&0&\;&0&\;&0&\;&\;\\ \;&0&\;&0&\;&0&\;&0&\;&0&\;\\ \mathllap{299}6&\;&\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\mathclap{20993}\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}&\;&\mathclap{15267}&\;&\mathclap{917}&\;&0&\;&0\\ \;&\mathllap{157}5&\;&0&\;&0&\;&0&\;&0&\;\\ \;&\;&\mathllap{577}5&\;&0&\;&0&\;&0&\;&\;\\ \;&\;&\;&\mathllap{1039}5&\;&0&\;&0&\;&\;&\;\\ \;&\;&\;&\;&\mathllap{900}2&\;&0&\;&\;&\;&\;\\ \;&\;&\;&\;&\;&\mathllap{299}6.&\;&\;&\;&\;&\;\\ \end{matrix}$
-And so we have by [Kü22, Theorem 4.14],
+And so we have by [Kü24, Theorem 4.14],
 $\dim\ker\left(f_{*}:\mathrm{HH}^{8}\left(X,\mathcal{O}_{X}\left(-8\right)\right)\to\mathrm{HH}^{8}\left(X,f_{*}\mathcal{O}_{X}\left(-8\right)\right)\right)$ $=\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}20993\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}.$
 Again every single $0\neq\eta$ in this $20993$-dimensional space give rise to a non-Fourier-Mukai functors.
-Finally, paired with [Kü22, Proposition 4.16] we get the following corollary
+Finally, paired with [Kü24, Proposition 4.16] we get the following corollary
 Corollary 7.17.